### v2

# Non-Linear Scoring Model for Translation Quality Evaluation

Abstract
Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words. However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition.
Building on the previously introduced Multi-Range framework [GladkoffEtAl2024], this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length. This approach presents empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale. We propose a two‑parameter model
$E(x)=a\ln\!\bigl(1+bx\bigr),\quad a,b>0,$
that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step. The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function.
It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations. By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation. The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation congruent with human judgment. Implementation in CAT/LQA systems and implications for human and AI-generated text evaluation are discussed.
1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision. It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014], with proper training of annotators. Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment.
In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample. This model typically normalizes errors per 1000 or 2000 words to compare them against a client-defined threshold, e.g., no more than five minor errors per 1000 words. The normalized error density is then used to make a pass-or-fail decision or compute a quality score.
This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words). However, problems arise when evaluating smaller or larger samples.
Clients, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality. Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability. This suggests that perceived translation quality is not a linear function. Evidence from other domains points in the same direction. In speech and video quality, international standards report top-end ’compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic]. Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ’showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450]. Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021]. And in usability engineering, a single ’catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994]. Together, these applied literatures reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
In points‑off situations, the linear rule makes long samples appear better than experts judge them to be by raising the acceptance threshold in direct proportion to length.
A recent quote from a large institutional client illustrates this point:
’Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate. Instead, we would prefer to fail the sample already at three or four major errors. This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.’
Another client confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words. However, in real-world operations, such precisely consistent evaluation sample sizes are rare. Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance. Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations.
This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length. The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases. We capture this pattern with a logarithmic dependency.
Earlier work [GladkoffEtAl2024] introduced a logarithmic scorecard. This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory); we then provide a simple, easily calibrated model for use in current evaluation tools.
2 Background
2.1 Analytic TQE and MQM
Analytic Translation Quality Evaluation (TQE) refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications. Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories.
The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling. Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the client’s preference.
Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict. The MQM framework is currently the most detailed and accurate approach to analytic translation-quality evaluation in professional practice. Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score. For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words.
Refinements to either component therefore directly advance the state of the art in translation-quality measurement.
2.2 Linear Scoring Models and Their Limits
The dominant approach to scoring in translation quality evaluation is the linear model. Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size (typically 1,000–2,000 words) so that assessments are comparable across jobs of different lengths. For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
Although this proportional scaling appears objective, it fails however to reflect how human content users actually perceive quality. Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]; however, this tolerance wanes as the document length grows further and the errors begin to compound. For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance. Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
  • micro-range (<250<250 words): For very short samples (under 250 words), traditional quality scoring methods are unreliable. As shown in [gladkoff2021], quality scores for such brief samples exhibit extremely wide confidence intervals—meaning the range of possible true scores is so large that the measurement is imprecise—making individual measurements meaningless for decision-making.
    This limitation is particularly important for automatic Quality Estimation (QE), especially with the practice of sentence- or segment-level scoring. Recent industry findings confirm that these fine-grained predictions do not align with how humans perceive quality. Bane[baneATA2025] highlighted the repeated failures of sentence- or segment-level QE in high-volume translation, where automated scores fail to identify all major, and even critical, errors, making the method unreliable. As discussed in [gladkoff2025substack], attempting to derive meaningful scores from very small samples or isolated sentences or segments simply does not work.
    Instead, for micro-range samples, Statistical Quality Control (SQC) methods offer a more statistically robust and risk-aware framework for interpreting quality signals, explicitly framing the producer’s and consumer’s risk associated with very limited observations.
  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a client-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another linear model is needed; a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
The present study concentrates on the entire span of statistically reliable measurements in both the meso-range and the macro-range and supplies both psychophysical and empirical evidence for adopting a logarithmic scoring model across these ranges. Here “statistically reliable” is meant in the technical sense—i.e., estimates with reasonably narrow confidence intervals —the usual sampling-uncertainty bands around an estimate that shrink as sample size grows. [^1] For intuition, a simple 95% Wald interval for a proportion is $p\pm 1.96\sqrt{p(1-p)/n}$ [gladkoff2021]. For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or its closely related Agresti–Coull “plus-four” approximation—provides markedly better coverage [AgrestiCoull1998]. Wilson 95% CI: $\frac{\hat{p}+\frac{z^{2}}{2n}\pm z\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},$ with z=1.96z\!=\!1.96; the Agresti–Coull “plus-four” uses $\tilde{p}=(x+\tfrac{z^{2}}{2})/(n+z^{2})$ and $\tilde{p}\pm z\sqrt{\tfrac{\tilde{p}(1-\tilde{p})}{n+z^{2}}}$ [AgrestiCoull1998].
When linear scoring is applied indiscriminately across all ranges, it produces systematic friction. Clients and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples. Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment. Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably. [^2] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
The persistent use of linear models—despite long-recognized shortcomings—stems largely from the absence of simple, theoretically grounded alternatives. This paper aims to fill that gap by presenting a non-linear scoring model and providing empirical and theoretical support for it, grounded in psychophysics and cognitive-load research.
3 Empirical Evidence of Non-Linearity
To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large enterprise clients. Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
3.1 Survey Design
The extended calibration questionnaire described in [GladkoffEtAl2024] was reused, asking survey respondents to state acceptable error counts for multiple evaluation sample sizes without performing linear extrapolation. This method again proved effective in eliciting a latent non-linear response curve.
A detailed calibration questionnaire was designed to capture Quality Managers’ intuition about acceptable error tolerance—operationalized as the maximum number of severity-weighted penalty points (APT) they would still pass—across a range of evaluation sample sizes.
Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work. Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds. To counteract this bias, the clients’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
The questionnaire asked the following questions:
  • Intuitively, how many minor errors would you accept in a 1-page sample?
  • Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages?
  • For a 3-page text, how many minor errors seem tolerable—just based on your experience?
  • Don’t calculate—just estimate: how many minor errors would you allow in 4 pages?
  • Based on gut feeling, how many minor errors would still be okay in a 5-page sample?
  • Without applying your formal model, what number of minor errors seems acceptable in 10 pages?
  • Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
Elicitation controls.
To reduce linear anchoring, page sizes were presented in randomized order without showing prior answers; two page sizes were repeated at the end to probe intra‑rater stability. Respondents were instructed explicitly not to prorate from known thresholds and to answer from professional judgment alone.
The questionnaire deliberately focused on minor errors for two key reasons. First, higher-severity errors (major and critical) are too infrequent for reliable statistical analysis. Second, as many studies note, their impact on quality perception is significantly different and more pronounced. This specific focus is practical, as standard industry practice allows the severity of any error to be expressed as an equivalent number of minor errors.
While clients used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
3.2 Results
The most striking finding was a clear non-linear pattern in error tolerance. Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size. Instead, the tolerance curve rose steeply at first and then gradually flattened. This pattern was consistent across all three clients, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
Figures 13 illustrate the raw responses. For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable. In one case, both minor and major error series were collected.
Figure 1: Client 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
Figure 2: Client 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
Figure 3: Client 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
None of the clients’ tolerance curves followed a straight line. Although each client had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
The respondents themselves were surprised by this result. When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ’acceptable’ was non-linear. They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate.
This phenomenon aligns with everyday experience. In very short samples, a single error can be jarring and dominate the user’s entire impression of quality. In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness.
This degradation impacts more than just a pleasant "reading experience"; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly. Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
Consistent with Fig. 7 in [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size. The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions.
From Figures 1–3 it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
Quantitative comparison of linear vs. logarithmic fits (technical note)
For readers who want the formal evidence behind Fig. 3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics. Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination). AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
$\mathrm{SSE}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\qquad\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
where $n$ is the number of points and $k$ the number of fitted parameters. As Table 1 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments. These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
In other words, the visual impression is also statistically decisive.
Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points.
To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes. This is what the values in Table 1 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better). Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC. This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can. In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
Logarithmic $a\ln(1+bx)$ $3.353$ $0.590$ $1.551$ $0.471$ $0.945$ 6.550-6.550 6.658-6.658
Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
Table 1: Goodness-of-fit statistics for the Client 1 minor-error dataset ($x$ in pages; $E$ in minor points). Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit. Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
The standard indicators reported in Table 1 show how well a curve follows the observed data points. They illustrate why the logarithmic model offers a markedly superior fit than the linear one. [^3] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better). RMSE: the average size of that distance (smaller means the model sits almost on top of the data). $R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor). AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better). Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance. When multiple elicited points are available, we recommend weighted least squares (App. B), down‑weighting sparse or less reliable page sizes. Using the delta method in App. B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.
4 Psychophysical and Cognitive Foundations
Empirical findings from these client interviews show that perceived translation quality is not a simple linear function of error count or text length. Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950]. This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
4.1 Weber–Fechner Law and Logarithmic Perception
Psychophysics has long recognized that sensation grows non-linearly with stimulus intensity. In his 1834 monograph on weight discrimination, Weber reported that just-noticeable difference (JND) is a constant fraction of the baseline stimulus: “The smallest perceptible increase of weight bears a fixed ratio to the weight already lifted[weber1834, p. 152]. Three decades later, Fechner integrated Weber’s proportionality into a mathematical model of subjective sensation, arguing that “if equal stimulus ratios yield equal sensation increments, then sensation must be the logarithm of stimulus magnitude[fechner1860, vol. 1, p. 55]. This logarithmic relation—now known as the Weber–Fechner law—was the first quantitative statement of a compressive perceptual scale; an authoritative English translation is provided in [fechner1966].
Intuition.
Applied to translation quality evaluation, the law would predict that the subjective penalty of an additional error is highest when the text is nearly error-free and decreases as errors accumulate. In other words, the first mistranslation in a short paragraph feels much more jarring than the fifteenth typo in a 20-page report.
Subsequent psychophysical research refined, but did not overturn, this insight. Stevens showed that many modalities follow a power function $S=k\,I^{n}$ with exponents $n<1$ for brightness, loudness, and heaviness [stevens1957]. Although $n=0$ is not itself logarithmic, the power-law family $I^{n}$ is asymptotically log-like:
$\lim_{n\to 0}\frac{I^{\,n}-1}{n}\;=\;\ln I.$
Over finite dynamic ranges, exponents in the $0.2\text{–}0.4$ band produce curves that closely approximate a logarithmic profile. Over typical stimulus ranges, power exponents $n\in[0.2,0.4]$ are empirically indistinguishable from a log on plots like ours, so a logarithmic tolerance is the parsimonious choice here. Comprehensive reviews such as [gescheider1997] and the signal detection synthesis of [green1966] confirm that compressive (sublinear) mapping is the norm in vision, audition and touch.
Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales. Treating translation errors as discrete ’stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
4.2 Cognitive Load and Accumulated Impact of Errors
Whereas psychophysics explains the decrease in sensitivity to additional stimuli, cognitive psychology accounts for how these stimuli collectively strain mental resources. The theory of limited working memory, formalized as Cognitive Load Theory (CLT), provides a parallel foundation for non-linear perception of translation quality.
CLT was introduced by Sweller in the late 1980s as a way to explain why instructional design must respect the cognitive limitations of learners [sweller1988]. It builds on earlier findings such as Miller’s famous observation that the number of discrete items a person can hold in short-term memory is limited to approximately seven, plus or minus two [miller1956]. As Sweller notes, “Our capacity to process novel information is extremely limited, and when this limit is exceeded, learning and comprehension suffer[sweller1988, p. 257].
In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources. One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity. This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality.
Importantly, this degradation does not occur linearly. Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency. Studies in instructional design and reading comprehension (e.g. [mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001]. This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content.
Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework. Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption. Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
4.3 Perception and Cognition Converge on Logarithmic Scoring
Together, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically. A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors. This synthesis underpins the non-linear scoring model proposed in the next section.
5 Proposed Non-Linear Scoring Model
This paper adopts the terminology introduced in [GladkoffEtAl2024], distinguishing between three classes of scoring models.
  • Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”). This uses the sample’s penalty rate and is not a calibrated/anchored score.
  • Calibrated Linear: This scoring model maps a client’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
  • Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve. Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
Given both the empirical evidence and the theoretical justifications presented, this paper proposes a more nuanced application of scoring models rather than a complete replacement of the linear approach. While the traditional linear model is often considered sufficient for the meso-range (texts up to 3,000 words), our findings indicate this is precisely where its limitations can be most acute. For example, as shown later in this paper, a linear model calibrated for a 1,000-word reference sample loses significant accuracy (over 20%) when applied to a 2,000-word sample.
Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in §5.4). Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated. 444‘Calibration’ for the raw score simply means choosing the client’s pass/fail cut-point on the raw penalty-rate scale. Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No calibration mapping to a 0–100 quality scale is involved. Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring. To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages.
By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture. Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system. [^5] Other scoring models may be devised for specific applications. Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here. Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear. For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations. Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$. Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450]. (One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper. They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely. A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope]. LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution. In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size. The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
5.1 Limitations of Linear Extrapolation
Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size. For example, if a client accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc. However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
As noted above, many institutional clients describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound. In practice, however, evaluators do not extend that allowance proportionally. When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors. In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length. What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole. This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
5.2 Mathematical Model
This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
$E(x)\;=\;a\,\ln(1+b\,x),\qquad a>0,\;b>0.$ (1)
Throughout, $\ln$ denotes the natural logarithm.
If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged. This unit invariance ensures figures shown in pages and formulas written in words are consistent. The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens. This curve shape effectively combines both theoretical expectations and observed data.
To clarify on the form of this function, the intercept-free form $E(x)=a\,\ln(1+bx)$ is deliberately used, which enforces $E(0)=0$ and allows curvature control via $b$, not the Excel spreadsheet “logarithmic trendline” $c+k\ln x$. The latter introduces a free intercept and does not pass through the origin, so it cannot be anchored to calibration points in the same way.
5.3 Anchoring and Calibration Procedures
This paper distinguishes between two steps in adapting a scoring model to operational use:
  • Anchoring defines the maximum acceptable error limit for a standard amount of text. This tolerance is expressed in penalty points based on a client’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
  • Calibration takes the absolute penalty point total and maps it onto a normalized, more easily interpretable 0-100 quality scale for use in dashboards or QA reports.
Feasibility (two‑point). Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
$r:=\frac{E_{1}}{E_{0}},\qquad\rho:=\frac{x_{1}}{x_{0}},\qquad g(b):=\frac{\ln(1+bx_{1})}{\ln(1+bx_{0})}.$
For $b>0$, $g$ is continuous and strictly monotone, with $\lim_{b\downarrow 0}g(b)=\rho$ and $\lim_{b\uparrow\infty}g(b)=1$. Hence a unique $b^{\star}>0$ exists if and only if
$\boxed{\ \min\{1,\rho\}\;<\;r\;<\;\max\{1,\rho\}\ }.$
Intuitively, the longer sample must permit more errors but sublinearly; the shorter must permit fewer errors but not in direct proportion to length.
The model can be easily calibrated using two known tolerance points. Suppose we have:
  • The primary anchor point: This is the main tolerance defined during the Anchoring step, such as a maximum of 5 errors ($E_{0}=5$) for a reference sample size of 1,000 words ($x_{0}=1000$).
  • A second, empirical data point: This point is derived from observing the client’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
Two‑point calibration (numerical).
Let $r=E_{1}/E_{0}$ and define $f(b)=\ln(1+bx_{1})-r\,\ln(1+bx_{0})$. Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$ by the intermediate value theorem. Near $b=0$, $f(b)\approx b(x_{1}-rx_{0})$, so the sign at the origin is $\mathrm{sgn}(x_{1}-rx_{0})$. As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$; if $r>1$ the limit is -$\infty$, and if $r<1$ it is $\infty$, guaranteeing a sign change and thus a unique root. Solve $f(b)=0$ with a robust one‑dimensional method (e.g., bisection/Brent); then set
$a^{\star}=\frac{E_{0}}{\ln(1+b^{\star}x_{0})}.$
See  Appendix A. Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
Illustrative values.
Feasibility.
Ensure $r=E_{1}/E_{0}$ lies strictly between $1$ and $\rho=x_{1}/x_{0}$, i.e., $\min\{1,\rho\}<r<\max\{1,\rho\}$, before solving for $b$.
Solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving the calibrated curve
$E(x)\;=\;3.688\,\ln\!\bigl(1+0.00288\,x\bigr).$
Fallback for multiple points. If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available—or if the two chosen points violate the feasibility condition—estimate $(a,b)$ by constrained least squares:
$\min_{a>0,\;b>0}\ \sum_{i=1}^{n}\Bigl[E_{i}-a\,\ln\!\bigl(1+b\,x_{i}\bigr)\Bigr]^{2}.$
Given any $b_{0}>0$, a fast 1‑D search over $b$ with the closed‑form $a(b)=\dfrac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$ works well (then refine both jointly if desired).
This approach allows the model to be tailored to specific content types, organizational standards, or language pairs while still operating under a common mathematical framework. For instance, the error tolerance can be precisely defined for each case, ranging from extremely strict for software UI text, to moderately strict for marketing copy and documentation, and significantly more lenient for internal communications or user-generated content.
For small $b$, $\ln(1+b\,x)\approx b\,x$, so the model reduces locally to a linear tolerance rule; this explains why a calibrated linear slope performs acceptably only in the vicinity of its reference sample size.
If $x$ is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $W\cdot b$; $a$ is unchanged.
Figure 4: Three interpolations. The spreadsheet curve $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
5.4 Near-Reference Fidelity Interval for Linear Approximation (±20%)
In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used for convenience:
$E_{\mathrm{lin}}(x)\;=\;E_{\log}(x_{\mathrm{ref}})\,\frac{x}{x_{\mathrm{ref}}},\qquad E_{\log}(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr).$
To quantify where this linear approximation is acceptable, we define the ±20% fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
$\left|\frac{E_{\mathrm{lin}}(x)}{E_{\log}(x)}-1\right|\leq 0.20.$
Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the two boundary points admit a closed form via the Derivation (sketch). Set $E_{\rm lin}(x)=(x/x_{\rm ref})\,E_{\rm log}(x_{\rm ref})$ and require $E_{\rm lin}(x)=(1\pm\varepsilon)E_{\rm log}(x)$. Writing $\alpha=\ln(1+bx_{\rm ref})/(bx_{\rm ref})$ and $y=1+bx$, we obtain $\alpha\ln y=(1\pm\varepsilon)(y-1)/y$. Rearranging gives $y\,e^{-\frac{1\pm\varepsilon}{\alpha}\,y}=e^{-\frac{1\pm\varepsilon}{\alpha}}$, so
$y=-\frac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\Bigl(-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\Bigr),$
which yields the closed form in the next equation. The branch $W_{-1}$ selects the solution near $x_{\rm ref}$.
Lambert $W$ function:
$x_{\pm}\;=\;\frac{-\,\dfrac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\!\left(\,-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\right)-1}{\,b\,}$,
where $x_{-}$ (with the “-” sign) gives the lower boundary (linear is $0.8\times$ the log curve) and $x_{+}$ (with the “++” sign) gives the upper boundary (linear is $1.2\times$ the log curve). [^6] The branch $W_{-1}$ is selected to obtain the solution near $x_{\mathrm{ref}}$.
Let’s look at the numerical intervals for the example calibration. Using the illustrative curve from §5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the ±20% fidelity intervals are:
  • $\boldsymbol{x_{\mathrm{ref}}=1000}$ words: $x\in[579,\;1460]$ words, width $\approx 881$ words (about $42\%$ to $46\%$ relative to $x_{\mathrm{ref}}$).
  • $\boldsymbol{x_{\mathrm{ref}}=2000}$ words: $x\in[1307,\;2747]$ words, width $\approx 1440$ words (about $34.7\%$ to $37.4\%$ relative to $x_{\mathrm{ref}}$).
These calculations illustrate how far one can move away from the reference sample size while still retaining acceptable measurement accuracy. The window is relatively narrow: if the reference is 1,000 words, then evaluating a 2,000-word sample already falls outside the ±20% fidelity band and will not be accurate; likewise, if the reference is 2,000 words, applying the same linear rule to a 1,000-word sample is equally unreliable.
In absolute terms, the ±20% window widens as the reference size increases (from about 0.88 k to about 1.44 k words), but in relative terms it becomes proportionally narrower. This quantifies the earlier point: a calibrated linear rule is only reliable near its anchoring point. Further away, it becomes too harsh for short samples and too lenient for long ones. When jobs regularly fall outside these fidelity intervals, the logarithmic model should be used directly, or the linear rule should be re-anchored to a more suitable reference size.
5.5 Visual comparison across three sample-size ranges
Figure 5: Linear model (blue) vs. human perception (PFT, red). Yellow bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule over‑ or under‑estimates true tolerance.
Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes. In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation wordcount).
The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes.
This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
Figure 6: Applicability across evaluation sample sizes. SQC (micro, $\lesssim\!250$ words), calibrated linear MQM (meso; reliable only within $\sim\!\pm 20\%$ of anchor), and non-linear MQM (macro). Vertical mark at $\sim\!250$ words flags the micro‑range where deterministic curves are unreliable. Y‑axis: error tolerance (penalty points).
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics. In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
For the meso- and macro-ranges, the choice of an appropriate scoring model depends not on the absolute length of any single text, but on the expected range of fluctuation in evaluation sample sizes. This paper’s findings point to a crucial rule of thumb: if evaluation sample sizes are expected to vary by more than approximately 20% from the anchor point, the non-linear model is required for accuracy. A calibrated linear model is only a viable tool when evaluation sample sizes remain consistently within this tight  20% band around the anchor. Counter-intuitively, because the logarithmic curve flattens at large values, a linear model can be more applicable in the macro-range than the meso-range, provided its anchor point is also in that higher range.
Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy. This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges.
Scenario Recommended model
EWC varies within $\pm 20\%$ around the anchor Calibrated linear model (re-anchor if the range shifts)
EWC routinely outside the $\pm 20\%$ band Use the non-linear logarithmic model directly
Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC); deterministic curves unreliable
Table 2: Operational guidance: which scoring model to use in practice.
Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
6 Practical Implications
To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry. Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception. Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations. Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (client-specific PT/MSV mapping).
6.1 Operational Guidelines for Quality Managers
One immediate benefit of the non-linear model is its ability to inform and standardize quality expectations across variable evaluation sample sizes. Quality managers can use the calibrated log function to define clear error tolerance thresholds for different content types or use cases, reducing ambiguity when evaluating translation outputs that differ in word count.
For example, a quality manager might specify: “Our reference wordcount is 2000 words with a tolerance of 7 minor errors. For evaluation sample sizes between 1500 and 3000 words, the standard linear approximation is acceptable. Outside that range, the logarithmic model must be used to adjust expectations accordingly.”
This shift allows quality managers and stakeholders to move beyond one-size-fits-all scoring rules and instead adopt quality scores that reflect actual cognitive effort and perceived severity. The result is more fair and interpretable evaluation outcomes, especially when comparing translation outputs across different content types or word counts.
7 Building your non-linear scorecard
Notation. The designations and symbols in this section follow the variable naming used in the ASTM working draft WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
7.1 Model and least-squares fit
We model the pass/fail tolerance as a function of sample size $x$ (words) by
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr),\qquad a>0,\;b>0.$ (2)
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, estimate $(a,b)$ by least squares
$\min_{a>0,b>0}\;\sum_{i=1}^{n}\bigl[E_{i}-a\,\ln(1+b\,x_{i})\bigr]^{2}.$
Profiling out $a$ yields, for any fixed $b>0$,
$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})},\qquad\text{then minimize }S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2}\text{ in }b>0,$ (3)
[…]
If $x$ was fitted in pages with $W$ words/page and you now compute in words, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$; $a$ is unchanged.
7.2 From tolerance to a calibrated score
For an evaluation of size $x=\text{EWC}$ (words), compute the allowed penalty points at the passing threshold
$E_{\text{allowed}}(x)\;=\;\hat{a}\,\ln\!\bigl(1+\hat{b}\,x\bigr).$ (4)
Let $\mathrm{APT}$ be the absolute penalty total from the MQM table. Define the quality fraction
$\mathrm{QF}(x):=1-\frac{\mathrm{APT}}{E_{\text{allowed}}(x)}.$ (5)
Map $\mathrm{QF}$ to the standard calibrated scale using Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
$\mathrm{OS}(x):=\mathrm{PT}+\mathrm{DPI}\,\mathrm{QF}(x),\qquad\mathrm{OS}_{\text{disp}}:=\min\{\mathrm{MSV},\,\max\{0,\mathrm{OS}(x)\}\}.$ (6)
The rating is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, else Fail.
Monotonicity and clipping. Since $\mathrm{OS}(x)=\mathrm{PT}+\mathrm{DPI}\bigl(1-\mathrm{APT}/E_{\mathrm{allowed}}(x)\bigr)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ grows; displaying $\mathrm{OS}_{\rm disp}=\min\{\mathrm{MSV},\max[0,\mathrm{OS}(x)]\}$ does not affect Pass/Fail, which always uses the un‑clipped inequality $\mathrm{APT}\leq E_{\mathrm{allowed}}(x)$.
7.3 Scorecard mapping
We now make the non-linear tolerance usable in everyday QA. The table below shows how to compute the scorecard fields from the evaluation word count (EWC), the fitted coefficients $\hat{a},\hat{b}$, and the severity-weighted penalty total (APT). The MQM error table itself is unchanged; the only change is that the allowed penalty $E_{\text{allowed}}(x)$ depends on the sample size.
Name Excel-style computation
Evaluation size $x$ (EWC) input
Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (see §5.3)
Allowed penalty at size $x$, $E_{\mathrm{allowed}}$ = a*LN(1 + b*EWC)
Quality fraction QFQF $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
Calibrated score $\mathrm{OS}$ = PT + (MSV-PT) * QF
Displayed score $\mathrm{OS}_{\mathrm{disp}}$ = MIN(MSV, MAX(0, OS))
Decision margin DMDM $E_{\mathrm{allowed}}-\mathrm{APT}$
Pass/Fail = IF(APT <= Eallowed,"PASS","FAIL")
Table 3: Scorecard mapping: variables and Excel-style computations.
This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line). The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual client tolerance function; the linear line is drawn for comparison only.
Same-curvature retargeting (optional). If you want a new anchor $E(x_{\text{ref}})=E_{0}$ while maintaining curvature, keep $b$ and solve
$a\;=\;\frac{E_{0}}{\ln(1+b\,x_{\text{ref}})}.$
This re-anchors the tolerance without changing the curve shape.
Micro-range caution. Below $\sim$250 words, deterministic tolerance curves are statistically unreliable; in this sample size range methods of Statistical Quality Control must be used rather than score curves. Quantitatively, if the underlying minor‑error rate is $p$ per word and we observe only $n$ words, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so the relative uncertainty scales as $1/\sqrt{np}$ and becomes huge when $np\approx 1$ (e.g., 1 expected error in 200 words at 5/1000).
Worked example (documentation content). Calibrated tolerance: $E(x)=3.688\,\ln(1+0.00288\,x)$. For $x=3000$ words, $E_{\mathrm{allowed}}(x)\approx 8.357$ points. If $\mathrm{APT}=7$, then $QF=1-7/8.357\approx 0.162$. With $\mathrm{PT}=80$, $\mathrm{MSV}=100$ ($\mathrm{DPI}=20$), $\mathrm{OS}\approx 83.25$ and $DM\approx 1.36$ (Pass). If $\mathrm{APT}=9$, then $QF\approx-0.077$, $\mathrm{OS}\approx 78.46$ and $DM\approx-0.64$ (Fail).
7.4 Implications for CAT Tools and LQA Automation
The change in scoring logic has clear consequences for software systems that implement translation quality evaluation, particularly CAT tools with integrated LQA modules or standalone dedicated LQA platforms.
In the traditional linear model, these tools calculate scores based on a fixed error density relative to a single reference word count. However, with a logarithmic tolerance function, the score calculation must dynamically adjust based on the actual sample size. This means LQA modules and dashboards need to:
  • Replace static pass/fail thresholds based on a single anchor point with dynamic tolerance functions, $E(x)$, that adjust based on the actual word count ($x$).
  • Visualize quality scores against a curved baseline that represents the tolerance function, not a simple flat line.
  • Provide evaluators with interactive feedback showing how close a sample’s error count is to its unique, size-adjusted pass/fail threshold.
This functionality is particularly valuable for machine translation, post-editing, and AI-generated content, as evaluation samples from these sources often vary widely in word count.
7.5 Improvement of a Unified Calibrated Quality Scale
One major challenge in TQE is ensuring that quality scores remain human-readable, interpretable, and comparable across different types of content, language pairs, and evaluation sample lengths. The linear calibrated model is a significant improvement over using raw scores with very small passing intervals, but even the linear calibrated model can distort comparability when applied to large or small samples.
With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, client-specific quality scale. Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable.
For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers. The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
This normalization also improves the validity of quality measurement results analytics across a portfolio of projects, enabling better tracking of supplier performance, trend analysis, and root-cause investigations.
7.6 Applications Beyond Human Translation
While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content. Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation.
The model based on non-linear error tolerance provides a fairer and more realistic baseline for evaluating content. It avoids the common pitfalls of unduly penalizing shorter segments with one mistake or overlooking the accumulated minor issues in long, highly fluent AI-generated texts.
As organizations adopt more automated content generation, the ability to evaluate quality in a human-aligned, perceptually realistic manner becomes essential. The proposed scoring model provides a foundation for doing so.
7.7 Document-Level Evaluation and Future Automation
Translation quality never exists in a vacuum: every segment lives within the cohesive fabric of a document. A translation sample is therefore not a roll of toilet paper to be inspected centimeter by centimeter; its quality must be judged in the context of the larger text. The larger the sample relative to the document, the more faithfully that judgment reflects holistic properties such as cohesion, coherence, register, and terminological consistency.
Small extracts can still provide a rapid signal, but offer only a limited view of document quality. They cannot capture discourse-level phenomena, and—because error density interacts non-linearly with length—linear metrics break down outside their narrow reference window.
The MQM-based Non-Linear Calibrated model solves this by remaining simple and linear in the meso-range yet extending smoothly into the macro-range. As long as the evaluation sample is sufficiently large relative to the document, as recommended by the Multi-Range Framework[GladkoffEtAl2024]—it yields scores that are statistically reliable and consistent with the judgment of experts at the document level.
Of course, analytic TQE becomes increasingly expensive at large evaluation sample sizes. No current AI system can yet perform a complete analytic evaluation, but future models may acquire the ability to conduct formal TQE (full analytic MQM evaluation) over long contexts (document-scale samples), reliably capturing local errors (segment-level) and holistic properties (cohesion, coherence, terminology/style consistency, discourse). The logarithmic model proposed here is expressly designed to serve as the computational backbone for such future automation, ensuring that automated scoring remains perceptually valid across the full spectrum of document lengths.
7.8 Summary
In practical terms, the transition to a non-linear scoring model means:
  • Greater alignment between future automated document-level scoring systems and expert human judgment,
  • Improved fairness and consistency across evaluation samples of different lengths,
  • More intelligent quality evaluation systems producing evaluations closer to human perception,
  • Stronger foundations for evaluating both human and AI-generated translation.
This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline.
The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary. It is simple to implement, easy to calibrate, and theoretically grounded. By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
8 Limitations and Future Work
The following limitations and research directions qualify the benefits summarized in §7.8 and the conclusion.
While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged. Addressing these will be essential for refining the model and validating its broader applicability.
8.1 Limited Survey Group Size
We conducted structured interviews with quality managers at three very large enterprise clients, and the results are corroborated by their internal practices. Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small. Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy. However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
Further research should involve a broader range of organizations—including LSPs, in-house localization teams, and public-sector translation departments—across different domains and language pairs. Ideally, future studies should also include controlled experiments where evaluators independently assess translation quality using both holistic judgment and formal scoring models, enabling direct comparison between model predictions and actual pass/fail outcomes.
8.2 Content-Type Specific Effects
Different content types (e.g., marketing, legal, technical, UI strings) may have different quality expectations and error tolerance. Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by client or use case.
We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters. However, this observation requires further validation through broader field data. A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend. Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms.
We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
8.3 Severity Weighting and Mixed Error Types
This paper concentrates on modeling tolerance in terms of penalty points, each of which typically represents one minor error. In practice, quality evaluations combine several error severities (minor, major, critical) penalized by specific multipliers and sometimes further scaled by different error-type weights applied for different content types.
Severity scaling (i.e., fixed severity weights such as minor/major/critical multipliers) is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points. This is separate from any error-type weights (e.g., terminology vs. grammar) a metric may also define. While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often. For these rarer severities, an integer-valued sequence may be preferable, as indicated by Client 3’s data in Fig. 3. That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
Future work should examine the form of such integer sequences, quantify how far they depart from a pure logarithmic curve, and determine the parameters that best describe severity scaling across different content types and languages.
8.4 Inter-Rater Reliability, IRR
When a linear model is used, the variability among human raters can be partly attributed to the fact that the model is only accurate at its anchoring point and loses precision elsewhere. In contrast, a calibrated non-linear model better reflects human judgment across a wide range of evaluation sample sizes. As a result, Inter-Rater Reliability (IRR) should increase significantly when using a non-linear calibrated scoring model, as it provides consistent and comparable measurements across different evaluation sample sizes, languages, and content types. More research is needed to quantify the improvement in inter-annotator agreement and to determine how much such a model reduces variance compared to traditional linear approaches in real-world annotation settings.
8.5 Extension to AI Evaluation Benchmarks
This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models. However, this claim has yet to be formally validated. AI benchmarks which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
Future work should involve testing the non-linear model as a normalization layer in LLM evaluation tasks, particularly those involving paragraph- or document-level generation, where holistic perception plays a major role.
A further implication is that a calibrated non-linear scoring model could eventually serve as a reference framework for future AI-based document-level evaluations, which currently provide holistic but qualitative judgments without a quantitative calibration.
9 Conclusion
Translation quality evaluation plays a critical role in ensuring that multilingual content meets the linguistic, technical, and functional standards required for global communication. While current industry practices predominantly rely on linear scoring models—extrapolating error thresholds proportionally from fixed reference wordcounts—this paper builds on empirical and practitioner evidence showing that linear models do not consistently align with human evaluation, motivating the adoption of non-linear alternatives.
This paper presents empirical data from experienced quality managers at large translation buyers, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size. This effect was evident across organizations, content types, and error categories and severities. This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance. This model is mathematically simple, empirically grounded, and easily adaptable to specific client requirements. It provides a flexible, yet rigorous alternative to traditional linear error density calculations.
Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
The implications of adopting this model are wide-ranging. Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths. CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics. Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output.
While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear. By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable. In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method. [^7] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes. SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output). In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making.
This study consolidates the multi-range proposal of [GladkoffEtAl2024] by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
Acknowledgments
The authors gratefully acknowledge the MQM Council and the ASTM WK46396 Working Group (MQM 2.0: Analytic Translation Quality Evaluation) for providing a collaborative forum in which earlier versions of the non-linear scoring model were discussed, refined, and incorporated into the broader MQM framework. In particular, the authors would like to thank Ingemar Strandvik (MQM Council) and Angelika Vasaa, Head of the Quality Coordination Unit at the European Parliament, for their insightful comments, institutional perspective, and many valuable discussions that helped shape the ideas presented in this paper. Views and opinions expressed are however their own of the authors and acknowledged contributors, and do not necessarily reflect those of the European Union or Parliament. Neither the European Union nor the Parliament’s authority can be held responsible for them.
The authors thank the Quality Managers of the participating organizations (referred to in this document as Clients C1–C3) for their generous contributions of time, insight, and real-world data. Their candid responses and openness to critical reflection were essential to this research. Gratitude is also expressed to the broader translation quality evaluation community for ongoing discussions that continue to refine our shared understanding of what quality means in multilingual communication.
Appendix A. Calibration from Two Points (Numerical)
Checking the model feasibility.
Given two tolerance points $(x_{0},E_{0})$, $(x_{1},E_{1})$ with $x_{0}\neq x_{1}$, $E_{0},E_{1}>0$, define $r=E_{1}/E_{0}$, $\rho=x_{1}/x_{0}$. A unique calibration with $a>0,b>0$ exists iff
$\min\{1,\rho\}<r<\max\{1,\rho\}.$
One‑dimensional root‑finding. Let $f(b)=\ln(1+bx_{1})-r\ln(1+bx_{0})$. Near the origin, $f(b)\approx b(x_{1}-rx_{0})$. As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$, which flips sign relative to the origin when the feasibility condition holds.
Bisection (robust).
  1. 1.
    Compute $r=E_{1}/E_{0}$ and verify feasibility.
  2. 2.
    Set $b_{\mathrm{lo}}\leftarrow 0^{+}$ and evaluate $f(b_{\mathrm{lo}})$.
  3. 3.
    Grow $b_{\mathrm{hi}}$ geometrically until $f(b_{\mathrm{lo}})\,f(b_{\mathrm{hi}})<0$.
  4. 4.
    Bisect until convergence: $b\leftarrow(b_{\mathrm{lo}}+b_{\mathrm{hi}})/2$ and update the bracket by the sign of $f(b)$.
  5. 5.
    Output $b^{\star}$ and $a^{\star}=E_{0}/\ln(1+b^{\star}x_{0})$.
Notes. (i) Changing the log base rescales $a$ only; we use the natural logarithm $\ln$. (ii) For $b\approx 0$, $\ln(1+bx)\approx bx$, so the model reduces locally to a linear rule, clarifying why a single linear slope works only near its reference sample size. (iii) With $n\geq 3$ points $\{(x_{i},E_{i})\}$, fit $(a,b)$ by nonlinear least squares under $a>0,b>0$.
Appendix B. Least-Squares Calibration from Multiple Points
Problem.
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ with $x_{i}>0$ and $E_{i}>0$, estimate $a>0,b>0$ in
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr)$
by (constrained) least squares:
$\min_{a>0,\;b>0}\;S(a,b)\;\;:=\;\;\sum_{i=1}^{n}\Bigl[E_{i}-a\,L_{i}(b)\Bigr]^{2},\qquad L_{i}(b):=\ln\!\bigl(1+b\,x_{i}\bigr).$
Profiling out $a$.
For any fixed $b>0$, the minimizer in $a$ is
$a(b)\;=\;\frac{\sum_{i}E_{i}\,L_{i}(b)}{\sum_{i}L_{i}^{2}(b)}.$
Substituting yields the one‑dimensional profiled objective
$S(b)\;=\;\sum_{i}E_{i}^{2}\;-\;\frac{\Bigl(\sum_{i}E_{i}\,L_{i}(b)\Bigr)^{\!2}}{\sum_{i}L_{i}^{2}(b)}\;=\;S_{0}-\frac{S_{1}(b)^{2}}{S_{2}(b)},$
where $S_{0}=\sum_{i}E_{i}^{2}$, $S_{1}(b)=\sum_{i}E_{i}\,L_{i}(b)$, and $S_{2}(b)=\sum_{i}L_{i}^{2}(b)$.
Derivative for 1‑D solvers (optional).
Let $L_{i}^{\prime}(b)=\dfrac{x_{i}}{1+b\,x_{i}}$. Then
$S_{1}^{\prime}(b)=\sum_{i}E_{i}\,L_{i}^{\prime}(b),\qquad S_{2}^{\prime}(b)=2\sum_{i}L_{i}(b)\,L_{i}^{\prime}(b),$
and
$S^{\prime}(b)\;=\;-\,\frac{2\,S_{1}(b)\,S_{1}^{\prime}(b)\,S_{2}(b)\;-\;S_{1}(b)^{2}\,S_{2}^{\prime}(b)}{S_{2}(b)^{2}}.$
While not strictly required, $S^{\prime}(b)$ enables Newton/Brent methods. In practice, $S(b)$ is unimodal over $b>0$ for the elicited data we observed.
Numerically stable 1‑D procedure.
  1. 1.
    If a feasible two‑point calibration is available ( Appendix A. Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
  2. 2.
    Define $q(b)=S(b)$. Bracket a minimum by expanding geometrically from $b_{0}$ (e.g., multiply by $2$) until $q$ increases on both sides.
  3. 3.
    Minimize $q(b)$ on the bracket with a derivative‑free method (Brent or golden‑section).
  4. 4.
[…]
Weighted least squares (optional).
If different points have different reliabilities, use positive weights $w_{i}$ to minimize
$S_{w}(a,b)=\sum_{i}w_{i}\bigl[E_{i}-a\,L_{i}(b)\bigr]^{2}$
with
$a_{w}(b)=\frac{\sum_{i}w_{i}E_{i}\,L_{i}(b)}{\sum_{i}w_{i}L_{i}^{2}(b)},\qquad S_{w}(b)=\sum_{i}w_{i}E_{i}^{2}-\frac{\bigl(\sum_{i}w_{i}E_{i}L_{i}(b)\bigr)^{2}}{\sum_{i}w_{i}L_{i}^{2}(b)}.$
Then proceed as above with $S_{w}(b)$.
Approximate standard errors.
[…]
$\frac{\partial r_{i}}{\partial a}=-L_{i}(\hat{b}),\qquad\frac{\partial r_{i}}{\partial b}=-\,\hat{a}\,L_{i}^{\prime}(\hat{b})\;=\;-\,\hat{a}\,\frac{x_{i}}{1+\hat{b}\,x_{i}}.$
[…]
$\widehat{\mathrm{Cov}}(\hat{a},\hat{b})\;=\;\hat{\sigma}^{2}\,(J^{\top}J)^{-1}.$
For any $x>0$, the delta method gives the variance of predicted tolerance
$\widehat{\mathrm{Var}}\!\left[E(x)\right]\;\approx\;\nabla_{\theta}E(x)^{\top}\,\widehat{\mathrm{Cov}}(\hat{a},\hat{b})\,\nabla_{\theta}E(x),\quad\nabla_{\theta}E(x)=\begin{bmatrix}\ln(1+\hat{b}\,x)\\[3.0pt] \hat{a}\,\dfrac{x}{1+\hat{b}\,x}\end{bmatrix}$.
A nonparametric alternative is a bootstrap over the elicited points.
Implementation notes.
(i) Use $\ln$ for the natural log and enforce $b>0$. (ii) Rescale $x$ if needed to keep $1+\hat{b}\,x$ well‑conditioned numerically. (iii) When points are very near two‑point feasible, the LS minimum occurs close to the two‑point solution.
[…]
Sensitivity to anchors (illustrative). With the second point fixed at $(x_{1},E_{1})=(250,2)$, varying the primary anchor by $\pm 1$ at $x_{0}=1000$ yields:
Anchor $E_{0}$ $E(1000)$ $E(2000)$ $E(3000)$
4 4.00 5.16 5.86
5 (baseline) 5.00 7.05 8.36
6 6.00 9.30 11.59
Near the anchor, predictions are stable by construction; farther away, two‑point calibrations diverge—hence the recommendation to use least squares with multiple points when available (this Appendix B).
Appendix C. Excel Goal Seek method for Two-Point Calibration
  1. 1.
    Put $x_{0}$ in A1, $E_{0}$ in B1, $x_{1}$ in C1, and $E_{1}$ in D1.
  2. 2.
    Enter an initial guess for $b>0$ in E1.
  3. 3.
    Compute $\varphi(b)$ in cell F1 using the formula:
    =LN(1 + \$E\$1 * \$C\$1) - (\$D\$1 / \$B\$1) * LN(1 + \$E\$1 * \$A\$1)
  4. 4.
    Go to Data $\rightarrow$ What-If Analysis $\rightarrow$ Goal Seek: set cell F1 to 0 by changing cell E1.
  5. 5.
    Once Goal Seek converges, compute $a$ in cell G1 using the formula:
    =\$B\$1 / LN(1 + \$E\$1 * \$A\$1)
Appendix D. Minimal Python code for Two-Point Calibration
Below is a minimal two-point calibration function that computes the non-linear tolerance parameters $a$ (overall scale) and $b$ (curvature) as described in the paper.
import math
def two_point_calibrate(x0, E0, x1, E1):
    r = E1 / E0
    def f(b):  # score-equation for b
        return math.log(1.0 + b*x1) - r*math.log(1.0 + b*x0)
    # bracket a root
    blo = 1e-12
    f_lo = f(blo)
    bhi = 1e-6
    for _ in range(80):
        f_hi = f(bhi)
        if f_lo * f_hi < 0:
            break
        bhi *= 2.0
    else:
        raise ValueError("Could not bracket a root for b; check inputs.")
    # bisection
    for _ in range(120):
        bmid = 0.5 * (blo + bhi)
        f_mid = f(bmid)
        if abs(f_mid) < 1e-12:
            blo = bhi = bmid
            break
        if f_lo * f_mid < 0:
            bhi = bmid
        else:
            blo = bmid
            f_lo = f_mid
    b = 0.5 * (blo + bhi)
    a = E0 / math.log(1.0 + b*x0)
    return a, b
Command-line usage: Save as calibrate.py and append the wrapper below to call it from a terminal:
if __name__ == "__main__":
    import argparse, math
    p = argparse.ArgumentParser(
        description="Two-point calibration for E(x)=a*ln(1+b*x)")
    p.add_argument("--x0", type=float, required=True, help="reference size x0")
    p.add_argument("--E0", type=float, required=True, help="tolerance at x0")
    p.add_argument("--x1", type=float, required=True, help="second size x1")
    p.add_argument("--E1", type=float, required=True, help="tolerance at x1")
    p.add_argument("--x",  type=float, help="optional size to evaluate E(x)")
    args = p.parse_args()
    a, b = two_point_calibrate(args.x0, args.E0, args.x1, args.E1)
    print(f"a={a:.6f}, b={b:.8g}")
    if args.x is not None:
        E = a * math.log(1.0 + b*args.x)
        print(f"E({args.x:g})={E:.6f}")
Example 1 (reproduce §5.3).
python3 calibrate.py --x0 1000 --E0 5 --x1 250 --E1 2
Expected output (approx.): a=3.688000, b=0.00288
Example 2 (compute E(3,000)E(3{,}000), §7.3).
python3 calibrate.py --x0 1000 --E0 5 --x1 250 --E1 2 --x 3000
Expected output (approx.): E(3000)=8.357
Appendix E. Least-Squares Fit (multiple points)
This routine fits the logarithmic tolerance $E(x)=a\,\ln(1+bx)$ to multiple tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ by profiling out $a$ and minimizing the one-dimensional objective in $b$ (see App. B). Given any $b>0$,
$a(b)=\frac{\sum_{i}E_{i}\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}\!,\quad S(b)=\sum_{i}\!\left(E_{i}-a(b)\ln(1+bx_{i})\right)^{2},$
[…]
Python code for Least Squares Fit:
import math
def calibrate_log_lsq(points, weights=None, b0=None, grow=2.0, tol=1e-8):
    Least-squares fit of E(x) = a * ln(1 + b x) to multiple points.
    points  : list of (x, E) pairs with x>0, E>0
    weights : optional list of positive weights w_i (same length as points)
    b0      : optional initial guess for b; default 1 / (10 * max x)
    returns : (a_hat, b_hat, sse)
    xs, Es = zip(*points)
    n = len(xs)
    if weights is None:
        ws = [1.0] * n
    else:
        if len(weights) != n:
            raise ValueError("weights must match points length")
        ws = list(weights)
    xmax = max(xs)
    if b0 is None or b0 <= 0:
        b0 = 1.0 / (10.0 * xmax)
    def L(b):
        # vector of ln(1 + b x_i); return None if invalid
        try:
            return [math.log(1.0 + b * x) for x in xs]
        except ValueError:
            return None
    def a_of_b(b):
        Li = L(b)
        if Li is None or min(1.0 + b * x for x in xs) <= 0:
            return None
        num = sum(w * E * l for w, E, l in zip(ws, Es, Li))
        den = sum(w * l * l for w, l in zip(ws, Li))
        if den <= 0:
            return None
        return num / den
    def sse_profile(b):
        a = a_of_b(b)
        if a is None:
            return float("inf")
        Li = L(b)
        r = [math.sqrt(w) * (E - a * l) for w, E, l in zip(ws, Es, Li)]
        return sum(ri * ri for ri in r)
    # ---- bracket a minimum around b0 by geometric expansion
    a_l = b0 / grow
    a_c = b0
    a_r = b0 * grow
    f_l = sse_profile(a_l)
    f_c = sse_profile(a_c)
    f_r = sse_profile(a_r)
    # ensure f_c is the smallest
    it = 0
    while not (f_c <= f_l and f_c <= f_r) and it < 80:
        it += 1
        if f_l < f_r:
            a_r, f_r = a_c, f_c
            a_c, f_c = a_l, f_l
            a_l /= grow
            f_l = sse_profile(a_l)
        else:
            a_l, f_l = a_c, f_c
            a_c, f_c = a_r, f_r
            a_r *= grow
            f_r = sse_profile(a_r)
    if it == 80:
        raise ValueError("Could not bracket a minimum for b; check data.")
    # ---- golden-section search on [a_l, a_r]
    phi = (math.sqrt(5) - 1) / 2
    left, right = a_l, a_r
    x1 = right - phi * (right - left)
    x2 = left + phi * (right - left)
    f1 = sse_profile(x1)
    f2 = sse_profile(x2)
    while (right - left) > tol * (abs(left) + abs(right) + 1.0):
        if f1 > f2:
            left = x1
            x1, f1 = x2, f2
            x2 = left + phi * (right - left)
            f2 = sse_profile(x2)
        else:
            right = x2
            x2, f2 = x1, f1
            x1 = right - phi * (right - left)
            f1 = sse_profile(x1)
    b_hat = 0.5 * (left + right)
    a_hat = a_of_b(b_hat)
    sse = sse_profile(b_hat)
    return a_hat, b_hat, sse
To test this code from the command line, add the wrapper below (or place it in a separate lsq_fit.py) to fit from a list of $(x,E)$ pairs and optionally evaluate $E(x)$ at a target size.
if __name__ == "__main__":
    import argparse, math, sys
    p = argparse.ArgumentParser(
        description="Least-squares fit for E(x)=a*ln(1+b*x)")
    p.add_argument("--xy", action="append", metavar="X,E",
                   help="data point as ’x,E’ (repeat for multiple points)")
    p.add_argument("--x", type=float, help="optional size to evaluate E(x)")
    args = p.parse_args()
    if not args.xy:
        sys.exit("Provide at least one --xy ’x,E’ pair.")
    pts = []
    for pair in args.xy:
        x_str, E_str = pair.split(",")
        pts.append((float(x_str), float(E_str)))
    a, b, sse = calibrate_log_lsq(pts)
    print(f"a={a:.6f}, b={b:.8g}, SSE={sse:.6f}")
    if args.x is not None:
        print(f"E({args.x:g})={a*math.log(1.0+b*args.x):.6f}")
Example 1 (Client 1, pages vs. minor errors; Fig. 3/Table 1).
python3 lsq_fit.py \
  --xy 2,2 --xy 3,3 --xy 4,4 --xy 5,5 --xy 7,6 --xy 10,7 --xy 20,8
[…]
Example 2 (evaluate the fitted curve at 12 pages).
python3 lsq_fit.py \
  --xy 2,2 --xy 3,3 --xy 4,4 --xy 5,5 --xy 7,6 --xy 10,7 --xy 20,8 \
  --x 12
[…]
Appendix F. Wolfram Notebook for Least Squares Fit
This appendix provides a Wolfram Notebook implementation of the least-squares calibration method described in Section 5.3 and detailed in Appendix A. The code can be pasted directly into a Wolfram Notebook and run cell by cell. It reproduces the profiled least-squares fitting procedure for the model
$E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr),\qquad a>0,\;b>0,$
using bracketing and golden-section search over $b$, with $a$ profiled in closed form.
Instructions
  1. 1.
    Open Wolfram Mathematica and create a new Notebook.
  2. 2.
    Copy the code blocks below into separate cells.
  3. 3.
    Evaluate the cells in order (Shift+Enter).
  4. 4.
    The final cell produces a plot and fitted formula.
Cell 1: Definitions
ClearAll[CalibrateLogLSQ];
Options[CalibrateLogLSQ] = {"Weights"->None,"InitialB"->Automatic,
   "BracketStep"->2.,"Tolerance"->1.*^-8};
CalibrateLogLSQ[data_List, opts:OptionsPattern[]] := Module[
  {x,E,n,wOpt,w,b0,step,tol,L,Lp,aOfB,sseProfile,bracketMin,
   goldenMin,br,bHat,aHat,yHat,resid,sse,sigma2,J,JTJ,cov},
  x=N[data[[All,1]]]; E=N[data[[All,2]]]; n=Length[x];
  If[AnyTrue[x,#<=0&]||AnyTrue[E,#<=0&],Return[\$Failed,Module]];
  wOpt=OptionValue["Weights"];
  w=If[wOpt===None,ConstantArray[1.,n],N[wOpt]];
  If[Length[w]=!=n||AnyTrue[w,#<=0&],Return[\$Failed,Module]];
  b0=Replace[OptionValue["InitialB"],Automatic->(1./(10. Max[x]))];
  step=OptionValue["BracketStep"]; tol=OptionValue["Tolerance"];
  L[b_?NumericQ]:=Log[1.+b*x];
  Lp[b_?NumericQ]:=x/(1.+b*x);
  aOfB[b_?NumericQ]:=Module[{Li=L[b],num,den},
    If[Min[1.+b*x]<=0,Return[Indeterminate]];
    num=Total[w*E*Li]; den=Total[w*Li*Li];
    If[den<=0||!NumericQ[den],Indeterminate,num/den]];
  sseProfile[b_?NumericQ]:=Module[{a=aOfB[b],Li,r},
    If[!NumericQ[a],Return[Infinity]];
    Li=L[b]; r=Sqrt[w]*(E-a*Li); r.r];
  bracketMin[f_,bStart_,s_:2.,maxIter_:60]:=Module[{a,b,c,fa,fb,fc,it=0},
    a=bStart/s; b=bStart; c=bStart*s; fa=f[a]; fb=f[b]; fc=f[c];
    While[!(fb<fa&&fb<fc)&&it<maxIter,
      it++;
      If[fa<fb,
        c=b;fc=fb; b=a;fb=fa; a=a/s; If[a<=0,a=b/(2 s)]; fa=f[a],
        a=b;fa=fb; b=c;fb=fc; c=c*s; fc=f[c]]];
    {a,b,c}];
  goldenMin[f_,a0_,c0_,t_:1.*^-8,maxIter_:200]:=Module[
    {phi=(Sqrt[5]-1)/2.,a=a0,c=c0,x1,x2,f1,f2,it=0},
    x1=c-phi(c-a); x2=a+phi(c-a); f1=f[x1]; f2=f[x2];
    While[(c-a)>t(Abs[a]+Abs[c]+1.)&&it<maxIter,
      it++;
      If[f1>f2, a=x1; x1=x2; f1=f2; x2=a+phi(c-a); f2=f[x2],
                 c=x2; x2=x1; f2=f1; x1=c-phi(c-a); f1=f[x1]]];
    .5(a+c)];
  br=bracketMin[sseProfile,b0,step];
  bHat=goldenMin[sseProfile,br[[1]],br[[3]],tol];
  aHat=aOfB[bHat]; If[!NumericQ[aHat],Return[\$Failed,Module]];
  yHat=aHat*L[bHat]; resid=E-yHat;
  sse=(Sqrt[w]*resid).(Sqrt[w]*resid);
  sigma2=sse/Max[n-2,1];
  J=Transpose[{-L[bHat],-aHat*Lp[bHat]}];
  JTJ=If[wOpt===None,Transpose[J].J,Transpose[Sqrt[w]*J].(Sqrt[w]*J)];
  cov=sigma2*PseudoInverse[JTJ];
  <|"a"->aHat,"b"->bHat,"SSE"->sse,"Sigma2"->sigma2,
    "Residuals"->resid,"Covariance"->cov,
    "Predict"->(aHat*Log[1.+bHat*#]&),"Bracket"->br[[{1,3}]]|>
Cell 2: Fit and Quick Results
data = {{2,2},{3,3},{4,4},{5,5},{7,6},{10,7},{20,8}};
fit  = CalibrateLogLSQ[data];
{fit["a"], fit["b"], fit["SSE"]} // N
Cell 3: Plot
 With[{a = fit["a"], b = fit["b"]},
 Show[
     ListPlot[data, PlotStyle -> {Blue, PointSize[.02]}],
     ListLinePlot[data, InterpolationOrder -> 1,
                  PlotStyle -> {Blue, Dotted, Thick}],
     Plot[a*Log[1 + b*x], {x, 0.5, 25}, PlotStyle -> {Red, Thick}]
   GridLines -> Automatic, Frame -> True,
   FrameLabel -> {"Pages (x)", "Tolerance  E(x)"},
   PlotLabel -> Style["Least-squares log fit", 14, Bold],
  Epilog -> {
  Inset[
    Style[
      Row[{"E(x) \[TildeTilde] ", NumberForm[a, {5,5}], " ln(1 + ",
           NumberForm[b, {5,5}], " x)"}],
      12, Bold, Red],
    Scaled[{0.6, 0.65}]
   ImageSize -> Large
Illustration and Resulting Formula
Figure 8 shows the illustration produced by Cell 3. The fitted formula is returned by {fit["a"], fit["b"]} in Cell 2, which evaluates numerically to
$E(x)\;\approx\;\hat{a}\,\ln\!\bigl(1+\hat{b}\,x\bigr),$
with coefficients $\hat{a},\hat{b}$ depending on the dataset.
For the sample dataset serving as example, the approximated function is:
$E(x)\;\approx\;3.35301\,\ln\!\bigl(1+0.59046\,x\bigr),$
with residual sum of squares (SSE) $\approx 1.55087$.
For different dataset change data array in Cell 2.
Figure 8: Least-squares logarithmic fit in Wolfram Notebook.

### v3

@@
-Non-Linear Scoring Model for Translation Quality Evaluation
+
 Abstract
@@
 However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition.
-Building on the previously introduced Multi-Range framework [GladkoffEtAl2024], this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length.
-This approach presents empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
+Building on our previously introduced Multi-Range framework, this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length.
+We present empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
 This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale.
-We propose a two‑parameter model
-$E(x)=a\ln\!\bigl(1+bx\bigr),\quad a,b>0,$
-that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step.
+We propose a two‑parameter model that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step.
 The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function.
@@
 By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation.
-The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation congruent with human judgment.
-Implementation in CAT/LQA systems and implications for human and AI-generated text evaluation are discussed.
+The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation, congruent with human judgment.
+keywords:
+Translation Quality Evaluation, Statistical Quality Measurement, Translation Evaluation Model, Non-Linear Scoring, Empirical-Based Design
 1 Introduction
 Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision.
-It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014], with proper training of annotators.
+It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], with proper training of annotators.
 Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment.
 In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample.
-This model typically normalizes errors per 1000 or 2000 words to compare them against a client-defined threshold, e.g., no more than five minor errors per 1000 words.
-The normalized error density is then used to make a pass-or-fail decision or compute a quality score.
+This model typically normalizes errors per 1000 or 2000 words to compare them against a user-defined threshold, e.g., no more than five minor errors per 1000 words.
+The normalized error density is then used to make a pass-or-fail decision or compute a quality score  [^1] Logrus Global Statistics https://logrusglobal.substack.com/archive .
 This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words).
-However, problems arise when evaluating smaller or larger samples.
-Clients, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality.
+However, problems arise when evaluating smaller or larger samples [gladkoff2021].
+Industrial users, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality.
 Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability.
@@
 Evidence from other domains points in the same direction.
-In speech and video quality, international standards report top-end ’compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic].
-Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ’showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450].
+In speech and video quality, international standards report top-end ‘compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic].
+Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ‘showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450].
 Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021].
-And in usability engineering, a single ’catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994].
-Together, these applied literatures reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
-In points‑off situations, the linear rule makes long samples appear better than experts judge them to be by raising the acceptance threshold in direct proportion to length.
-A recent quote from a large institutional client illustrates this point:
-’Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them.
-We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient.
-The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula.
-For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate.
-Instead, we would prefer to fail the sample already at three or four major errors.
-This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.’
-Another client confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words.
-However, in real-world operations, such precisely consistent evaluation sample sizes are rare.
-Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
+In usability engineering, a single ‘catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994].
+Together, these applied literature reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
 Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance.
 Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations.
-This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length.
+[^2] We list some real-world institutional users’ experiences on translation evaluation in Appendix A.
+This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in our earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length.
 The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases.
 We capture this pattern with a logarithmic dependency.
-Earlier work [GladkoffEtAl2024] introduced a logarithmic scorecard.
-This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory); we then provide a simple, easily calibrated model for use in current evaluation tools.
+The Multi-Range TQE work introduced a logarithmic scorecard.
+This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory).
+We also provide a simple, easily calibrated model for use in current evaluation tools.
+The structure of this paper is organized as below: Section 2 introduces the background on analytical TQE and MQM, Section 3 presents our empirical evidence of non-linearity, Section 4 offers psychophysical and cognitive foundation, Section 5 details the non-linear scoring model, Section 6 and 7 cover the practical implications and step-by-step guide on building a non-linear scorecard, followed by Limitations and Future Work (Section 8).
 2 Background
 2.1 Analytic TQE and MQM
-Analytic Translation Quality Evaluation (TQE) refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications.
+Analytic TQE refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications.
 Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories.
@@
 MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling.
-Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the client’s preference.
+Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the user’s preference.
 Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict.
-The MQM framework is currently the most detailed and accurate approach to analytic translation-quality evaluation in professional practice.
 Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score.
-For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words.
-Refinements to either component therefore directly advance the state of the art in translation-quality measurement.
+For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words  [^3] https://themqm.org/.
 2.2 Linear Scoring Models and Their Limits
 The dominant approach to scoring in translation quality evaluation is the linear model.
-Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size (typically 1,000–2,000 words) so that assessments are comparable across jobs of different lengths.
+Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size so that assessments are comparable across jobs of different lengths.
 For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
-Although this proportional scaling appears objective, it fails however to reflect how human content users actually perceive quality.
-Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]; however, this tolerance wanes as the document length grows further and the errors begin to compound.
+Although this proportional scaling appears objective, it fails to reflect how human content users actually perceive quality.
+Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450].
+However, this tolerance wanes as the document length grows further and the errors begin to compound.
 For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance.
 Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
-Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
+Building on our earlier work on Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
   • micro-range (<250<250 words): For very short samples (under 250 words), traditional quality scoring methods are unreliable.
@@
     Instead, for micro-range samples, Statistical Quality Control (SQC) methods offer a more statistically robust and risk-aware framework for interpreting quality signals, explicitly framing the producer’s and consumer’s risk associated with very limited observations.
-  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a client-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
-  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another linear model is needed; a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
+  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a user-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
+  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another model is needed: a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
 The present study concentrates on the entire span of statistically reliable measurements in both the meso-range and the macro-range and supplies both psychophysical and empirical evidence for adopting a logarithmic scoring model across these ranges.
@@
 When linear scoring is applied indiscriminately across all ranges, it produces systematic friction.
-Clients and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples.
+Users and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples.
 Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment.
@@
 3 Empirical Evidence of Non-Linearity
-To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large enterprise clients.
+To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large institutional users (anonymized).
 Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
@@
 Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds.
-To counteract this bias, the clients’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
+To counteract this bias, the institutions’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
 The questionnaire asked the following questions:
-  • Intuitively, how many minor errors would you accept in a 1-page sample?
+Intuitive Error Tolerance Questions Intuitively, how many minor errors would you accept in a 1-page sample?
 Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages?
 For a 3-page text, how many minor errors seem tolerable—just based on your experience?
-  • Don’t calculate—just estimate: how many minor errors would you allow in 4 pages?
+Do not calculate—just estimate: how many minor errors would you allow in 4 pages?
 Based on gut feeling, how many minor errors would still be okay in a 5-page sample?
@@
 This specific focus is practical, as standard industry practice allows the severity of any error to be expressed as an equivalent number of minor errors.
-While clients used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
+While institutions used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
 3.2 Results
@@
 Instead, the tolerance curve rose steeply at first and then gradually flattened.
-This pattern was consistent across all three clients, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
+This pattern was consistent across all three anonymized institutions, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
 Figures 13 illustrate the raw responses.
@@
 In one case, both minor and major error series were collected.
-Figure 1: Client 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
-Figure 2: Client 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
-Figure 3: Client 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
-None of the clients’ tolerance curves followed a straight line.
-Although each client had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
+Figure 1: Institution 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
+Figure 2: Institution 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
+Figure 3: Institution 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
+None of the institutions’ tolerance curves followed a straight line.
+Although each organization had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
 The respondents themselves were surprised by this result.
-When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ’acceptable’ was non-linear.
+When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ‘acceptable’ was non-linear.
 They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate.
@@
 In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness.
-This degradation impacts more than just a pleasant "reading experience"; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly.
+This degradation impacts more than just a pleasant ”reading experience”; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly.
 Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
 Consistent with Fig.
-7 in [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
+7 in our pilot study from [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
 Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size.
 The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions.
-From Figures 1–3 it is already clear that the data points do not fall on a straight line. The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
-Quantitative comparison of linear vs. logarithmic fits (technical note)
+From Figures 13, it is already clear that the data points do not fall on a straight line.
+The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
 For readers who want the formal evidence behind Fig.
-3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics.
-Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination).
-AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
-$\mathrm{SSE}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\qquad\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
-$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
-where $n$ is the number of points and $k$ the number of fitted parameters.
-As Table 1 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments.
-These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
-In other words, the visual impression is also statistically decisive.
-Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points.
-To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes.
-This is what the values in Table 1 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better).
-Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC.
-This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can.
-In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
-Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
-Logarithmic $a\ln(1+bx)$ $3.353$ $0.590$ $1.551$ $0.471$ $0.945$ 6.550-6.550 6.658-6.658
-Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
-AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
-Table 1: Goodness-of-fit statistics for the Client 1 minor-error dataset ($x$ in pages; $E$ in minor points).
-Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit.
-Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
-The standard indicators reported in Table 1 show how well a curve follows the observed data points.
-They illustrate why the logarithmic model offers a markedly superior fit than the linear one.
-[^3] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better).
-RMSE: the average size of that distance (smaller means the model sits almost on top of the data).
-$R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor).
-AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better).
-Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance.
-When multiple elicited points are available, we recommend weighted least squares (App.
-B), down‑weighting sparse or less reliable page sizes.
-Using the delta method in App.
-B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.
+3: we detail the linear vs logarithmic fits in Appendix B.
 4 Psychophysical and Cognitive Foundations
-Empirical findings from these client interviews show that perceived translation quality is not a simple linear function of error count or text length.
+Empirical findings from these anonymized institution interviews show that perceived translation quality is not a simple linear function of error count or text length.
 Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950].
@@
 Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales.
-Treating translation errors as discrete ’stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
+Treating translation errors as discrete ‘stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
 4.2 Cognitive Load and Accumulated Impact of Errors
@@
 4.3 Perception and Cognition Converge on Logarithmic Scoring
-Together, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically.
+In summary, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically.
 A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors.
@@
 5 Proposed Non-Linear Scoring Model
-This paper adopts the terminology introduced in [GladkoffEtAl2024], distinguishing between three classes of scoring models.
+This paper adopts the terminology introduced in or primary study [GladkoffEtAl2024], distinguishing between three classes of scoring models.
 Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”).
 This uses the sample’s penalty rate and is not a calibrated/anchored score.
-  • Calibrated Linear: This scoring model maps a client’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
+Calibrated Linear: This scoring model maps a user/institution’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
 Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve.
@@
 For example, as shown later in this paper, a linear model calibrated for a 1,000-word reference sample loses significant accuracy (over 20%) when applied to a 2,000-word sample.
-Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in §5.4).
+Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4).
 Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated.
-444‘Calibration’ for the raw score simply means choosing the client’s pass/fail cut-point on the raw penalty-rate scale.
+666‘Calibration’ for the raw score simply means choosing the user’s pass/fail cut-point on the raw penalty-rate scale.
 Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words).
@@
 Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system.
-[^5] Other scoring models may be devised for specific applications.
+5.1 Review the Limitations of Linear Extrapolation
+Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size.
+For example, if a user accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc.
+However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
+As noted above, surveyed institutional users describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound.
+In practice, however, evaluators do not extend that allowance proportionally.
+When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors.
+In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length.
+What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole.
+This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
+5.2 Proposed Mathematical Model of Acceptable Errors
+This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
+$E(x)\;=\;a\,\ln(1+b\,x),\qquad a>0,\;b>0.$ (1)
+Throughout, $\ln$ denotes the natural logarithm.
+If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged.
+This unit invariance ensures figures shown in pages and formulas written in words are consistent.
+The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens.
+This curve shape effectively combines both theoretical expectations and observed data.
+To clarify the functional form, we adopt the intercept-free specification $E(x)=a\,\ln(1+bx)$.
+This choice enforces the natural boundary condition $E(0)=0$ and provides explicit control over curvature through the parameter $b$.
+This formulation is distinct from the “logarithmic trendline” $c+k\ln x$ commonly implemented in spreadsheet software, which introduces a free intercept term and does not pass through the origin.
+Consequently, the latter cannot be anchored to calibration points in an equivalent manner.
+[^7] A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope].
+LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution.
+In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size.
+The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
+Other scoring models may be devised for specific applications.
 Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here.
@@
 They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely.
-A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope].
-LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution.
-In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size.
-The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
-5.1 Limitations of Linear Extrapolation
-Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size.
-For example, if a client accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc.
-However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
-As noted above, many institutional clients describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound.
-In practice, however, evaluators do not extend that allowance proportionally.
-When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors.
-In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length.
-What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole.
-This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
-5.2 Mathematical Model
-This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
-$E(x)\;=\;a\,\ln(1+b\,x),\qquad a>0,\;b>0.$ (1)
-Throughout, $\ln$ denotes the natural logarithm.
-If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged.
-This unit invariance ensures figures shown in pages and formulas written in words are consistent.
-The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens.
-This curve shape effectively combines both theoretical expectations and observed data.
-To clarify on the form of this function, the intercept-free form $E(x)=a\,\ln(1+bx)$ is deliberately used, which enforces $E(0)=0$ and allows curvature control via $b$, not the Excel spreadsheet “logarithmic trendline” $c+k\ln x$.
-The latter introduces a free intercept and does not pass through the origin, so it cannot be anchored to calibration points in the same way.
 5.3 Anchoring and Calibration Procedures
@@
   • Anchoring defines the maximum acceptable error limit for a standard amount of text.
-This tolerance is expressed in penalty points based on a client’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
+This tolerance is expressed in penalty points based on a user’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
   • Calibration takes the absolute penalty point total and maps it onto a normalized, more easily interpretable 0-100 quality scale for use in dashboards or QA reports.
-Feasibility (two‑point). Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
+Feasibility (two‑point).
+Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
 $r:=\frac{E_{1}}{E_{0}},\qquad\rho:=\frac{x_{1}}{x_{0}},\qquad g(b):=\frac{\ln(1+bx_{1})}{\ln(1+bx_{0})}.$
@@
   • The primary anchor point: This is the main tolerance defined during the Anchoring step, such as a maximum of 5 errors ($E_{0}=5$) for a reference sample size of 1,000 words ($x_{0}=1000$).
-  • A second, empirical data point: This point is derived from observing the client’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
+  • A second, empirical data point: This point is derived from observing the user’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
 Two‑point calibration (numerical).
@@
 $a^{\star}=\frac{E_{0}}{\ln(1+b^{\star}x_{0})}.$
-See  Appendix A. Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
+See  Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
 Illustrative values.
@@
 Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes.
-In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation wordcount).
+In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation word-count).
 The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes.
@@
 Y‑axis: error tolerance (penalty points).
-Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics.
+In summary, building on the Multi-Range Framework [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics.
 In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
@@
 This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges.
+Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
 Scenario Recommended model
@@
 Table 1: Operational guidance: which scoring model to use in practice.
-Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
 6 Practical Implications
@@
 Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations.
-Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (client-specific PT/MSV mapping).
+Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (user-specific PT/MSV mapping).
 6.1 Operational Guidelines for Quality Managers
@@
 7 Building your non-linear scorecard
-Notation.
-The designations and symbols in this section follow the variable naming used in the ASTM working draft WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
+We list the following steps in this section including model and least-squares fit, from tolerance to a calibrated score, and scorecard mapping, followed by implications for CAT tools and linguistic quality assessment (LQA) automation, improvement of a unified calibrated quality scale, applications beyond human translation, and document-level evaluation.
+[^9] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
+http://www.astm.org/workitem-wk46396
 7.1 Model and least-squares fit
@@
 Table 2: Scorecard mapping: variables and Excel-style computations.
-This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
+This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via Equation  (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
 Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line).
-The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual client tolerance function; the linear line is drawn for comparison only.
+The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual user tolerance function; the linear line is drawn for comparison only.
 Same-curvature retargeting (optional). If you want a new anchor $E(x_{\text{ref}})=E_{0}$ while maintaining curvature, keep $b$ and solve
@@
 The linear calibrated model is a significant improvement over using raw scores with very small passing intervals, but even the linear calibrated model can distort comparability when applied to large or small samples.
-With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, client-specific quality scale.
+With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, user-specific quality scale.
 Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable.
@@
 While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content.
-Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation.
+Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation [han2024neural].
 The model based on non-linear error tolerance provides a fairer and more realistic baseline for evaluating content.
@@
 8 Limitations and Future Work
-The following limitations and research directions qualify the benefits summarized in §7.8 and the conclusion.
+The following limitations and research directions qualify the benefits summarized in Section 7.8 and the conclusion.
 While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged.
@@
 8.1 Limited Survey Group Size
-We conducted structured interviews with quality managers at three very large enterprise clients, and the results are corroborated by their internal practices.
+We conducted structured interviews with quality managers at three very large enterprise institutes, and the results are corroborated by their internal practices.
 Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small.
@@
 8.2 Content-Type Specific Effects
-Different content types (e.g., marketing, legal, technical, UI strings) may have different quality expectations and error tolerance.
-Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by client or use case.
+Different content types such as marketing, legal, technical, and UI strings, may have different quality expectations and error tolerance.
+Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by user or use case.
 We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters.
@@
 In practice, quality evaluations combine several error severities (minor, major, critical) penalized by specific multipliers and sometimes further scaled by different error-type weights applied for different content types.
-Severity scaling (i.e., fixed severity weights such as minor/major/critical multipliers) is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points.
+Severity scaling, i.e., fixed severity weights such as minor/major/critical multipliers, is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points.
 This is separate from any error-type weights (e.g., terminology vs.
@@
 While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often.
-For these rarer severities, an integer-valued sequence may be preferable, as indicated by Client 3’s data in Fig.
+For these rarer severities, an integer-valued sequence may be preferable, as indicated by Institution 3’s data in Fig.
 3.
@@
 However, this claim has yet to be formally validated.
-AI benchmarks which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
+AI benchmarks such as WMT [^10] https://machinetranslate.org/wmt shared task series [kocmi-etal-2025-findings] which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
 Future work should involve testing the non-linear model as a normalization layer in LLM evaluation tasks, particularly those involving paragraph- or document-level generation, where holistic perception plays a major role.
@@
 While current industry practices predominantly rely on linear scoring models—extrapolating error thresholds proportionally from fixed reference wordcounts—this paper builds on empirical and practitioner evidence showing that linear models do not consistently align with human evaluation, motivating the adoption of non-linear alternatives.
-This paper presents empirical data from experienced quality managers at large translation buyers, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size.
+This paper presents empirical data from experienced quality managers at large institutional translation service users, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size.
 This effect was evident across organizations, content types, and error categories and severities.
@@
 To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance.
-This model is mathematically simple, empirically grounded, and easily adaptable to specific client requirements.
+This model is mathematically simple, empirically grounded, and easily adaptable to specific user requirements.
 It provides a flexible, yet rigorous alternative to traditional linear error density calculations.
@@
 In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making.
-This study consolidates the multi-range proposal of [GladkoffEtAl2024] by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
-Acknowledgments
+This study consolidates the multi-range theory findings by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
+10 acknowledgments
 The authors gratefully acknowledge the MQM Council and the ASTM WK46396 Working Group (MQM 2.0: Analytic Translation Quality Evaluation) for providing a collaborative forum in which earlier versions of the non-linear scoring model were discussed, refined, and incorporated into the broader MQM framework.
@@
 Neither the European Union nor the Parliament’s authority can be held responsible for them.
-The authors thank the Quality Managers of the participating organizations (referred to in this document as Clients C1–C3) for their generous contributions of time, insight, and real-world data.
+The authors thank the Quality Managers of the participating organizations (referred to in this document as users C1–C3) for their generous contributions of time, insight, and real-world data.
 Their candid responses and openness to critical reflection were essential to this research.
 Gratitude is also expressed to the broader translation quality evaluation community for ongoing discussions that continue to refine our shared understanding of what quality means in multilingual communication.
-Appendix A. Calibration from Two Points (Numerical)
+Appendix A Institution User Quotations on Evaluation
+In points‑off situations, the linear rule makes long samples appear better than experts judge them to be by raising the acceptance threshold in direct proportion to length.
+A recent quote from a large institutional user illustrates this point:
+“Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them.
+We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient.
+The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula.
+For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate.
+Instead, we would prefer to fail the sample already at three or four major errors.
+This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.”
+Another industrial user confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words.
+However, in real-world operations, such precisely consistent evaluation sample sizes are rare.
+Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
+Appendix B Linear vs Logarithmic Fits
+For readers who want the formal evidence behind Fig.
+3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics.
+Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination).
+AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
+$\mathrm{SSE}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\qquad\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
+$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
+where $n$ is the number of points and $k$ the number of fitted parameters.
+As Table 3 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments.
+These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
+In other words, the visual impression is also statistically decisive.
+Below is the quantitative comparison of linear vs.
+logarithmic fits on the elicited points.
+To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes.
+This is what the values in Table 3 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better).
+Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC.
+This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can.
+In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
+Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
+Logarithmic $a\ln(1+bx)$ $3.353$ $0.590$ $1.551$ $0.471$ $0.945$ 6.550-6.550 6.658-6.658
+Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
+AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
+Table 3: Goodness-of-fit statistics for the user 1 minor-error dataset ($x$ in pages; $E$ in minor points).
+Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit.
+Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
+The standard indicators reported in Table 3 show how well a curve follows the observed data points.
+They illustrate why the logarithmic model offers a markedly superior fit than the linear one.
+[^12] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better).
+RMSE: the average size of that distance (smaller means the model sits almost on top of the data).
+$R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor).
+AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better).
+Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance.
+When multiple elicited points are available, we recommend weighted least squares (App.
+B), down‑weighting sparse or less reliable page sizes.
+Using the delta method in App.
+B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.
+Appendix C Calibration from Two Points (Numerical)
 Checking the model feasibility.
@@
 (iii) With $n\geq 3$ points $\{(x_{i},E_{i})\}$, fit $(a,b)$ by nonlinear least squares under $a>0,b>0$.
-Appendix B. Least-Squares Calibration from Multiple Points
+Appendix D Least-Squares Calibration from Multiple Points
 Problem.
@@
   1. 1.
-    If a feasible two‑point calibration is available ( Appendix A. Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
+    If a feasible two‑point calibration is available ( Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
   2. 2.
@@
 Near the anchor, predictions are stable by construction; farther away, two‑point calibrations diverge—hence the recommendation to use least squares with multiple points when available (this Appendix B).
-Appendix C. Excel Goal Seek method for Two-Point Calibration
+Appendix E Excel Goal Seek method for Two-Point Calibration
   1. 1.
@@
     =\$B\$1 / LN(1 + \$E\$1 * \$A\$1)
-Appendix D. Minimal Python code for Two-Point Calibration
+Appendix F Minimal Python code for Two-Point Calibration
 Below is a minimal two-point calibration function that computes the non-linear tolerance parameters $a$ (overall scale) and $b$ (curvature) as described in the paper.
@@
 Expected output (approx.): E(3000)=8.357
-Appendix E. Least-Squares Fit (multiple points)
+Appendix G Least-Squares Fit (multiple points)
 This routine fits the logarithmic tolerance $E(x)=a\,\ln(1+bx)$ to multiple tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ by profiling out $a$ and minimizing the one-dimensional objective in $b$ (see App.
@@
         print(f"E({args.x:g})={a*math.log(1.0+b*args.x):.6f}")
-Example 1 (Client 1, pages vs. minor errors; Fig. 3/Table 1).
+Example 1 (user 1, pages vs. minor errors; Fig. 3/Table 1).
 python3 lsq_fit.py \
@@
 […]
-Appendix F. Wolfram Notebook for Least Squares Fit
+Appendix H Wolfram Notebook for Least Squares Fit
 This appendix provides a Wolfram Notebook implementation of the least-squares calibration method described in Section 5.3 and detailed in Appendix A.
@@
 using bracketing and golden-section search over $b$, with $a$ profiled in closed form.
-Instructions
+H.1 Instructions
   1. 1.
@@
    ImageSize -> Large
-Illustration and Resulting Formula
+H.2 Illustration and Resulting Formula
 Figure 8 shows the illustration produced by Cell 3. The fitted formula is returned by {fit["a"], fit["b"]} in Cell 2, which evaluates numerically to

### v4

@@
 Abstract
-Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words.
-However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition.
-Building on our previously introduced Multi-Range framework, this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length.
-We present empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size.
-This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale.
-We propose a two‑parameter model that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step.
-The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function.
-It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations.
-By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation.
-The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation, congruent with human judgment.
+Analytic Translation Quality Evaluation (TQE) based on Multidimensional Quality Metrics (MQM) typically applies a linear error-to-penalty mapping calibrated on reference samples of 1000100020002000 words.
+However, extrapolating this scale to other sample lengths introduces systematic bias, over-penalizing short texts and under-penalizing long ones, and often conflicting with expert judgments.
+Building on our Multi-Range framework [GladkoffEtAl2024], we propose a calibrated non-linear scoring model that better reflects perceived translation quality across variable-length samples.
+Empirical evidence from three large-scale production evaluation settings indicates that acceptable error counts grow logarithmically, rather than linearly, with sample size.
+This trend aligns with psychophysical and cognitive accounts (e.g., the Weber–Fechner law and Cognitive Load Theory), which suggest diminishing marginal perceptual impact of errors alongside increasing cumulative cognitive burden.
+We introduce a two-parameter tolerance model anchored to a reference point and calibrated from two additional tolerance points via a one-dimensional root-finding procedure.
+The resulting formulation provides a closed-form range in which the linear approximation remains within $\pm 20\%$ relative error, and integrates into existing MQM workflows with minimal changes by replacing constant tolerance with a length-dependent function.
+The proposed method improves interpretability, fairness, and inter-rater reliability across both human and AI-generated translations, supporting more scalable and perceptually grounded translation quality evaluation.
 keywords:
-Translation Quality Evaluation, Statistical Quality Measurement, Translation Evaluation Model, Non-Linear Scoring, Empirical-Based Design
+Translation quality evaluation, MQM, Human evaluation, Non-linear scoring, Inter-rater reliability
 1 Introduction
-Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision.
-It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], with proper training of annotators.
-Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment.
-In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample.
-This model typically normalizes errors per 1000 or 2000 words to compare them against a user-defined threshold, e.g., no more than five minor errors per 1000 words.
-The normalized error density is then used to make a pass-or-fail decision or compute a quality score  [^1] Logrus Global Statistics https://logrusglobal.substack.com/archive .
-This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words).
-However, problems arise when evaluating smaller or larger samples [gladkoff2021].
-Industrial users, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality.
-Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability.
-This suggests that perceived translation quality is not a linear function.
-Evidence from other domains points in the same direction.
-In speech and video quality, international standards report top-end ‘compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic].
-Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ‘showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450].
-Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021].
-In usability engineering, a single ‘catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994].
-Together, these applied literature reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
-Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance.
-Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations.
-[^2] We list some real-world institutional users’ experiences on translation evaluation in Appendix A.
-This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in our earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length.
-The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases.
-We capture this pattern with a logarithmic dependency.
-The Multi-Range TQE work introduced a logarithmic scorecard.
-This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory).
-We also provide a simple, easily calibrated model for use in current evaluation tools.
-The structure of this paper is organized as below: Section 2 introduces the background on analytical TQE and MQM, Section 3 presents our empirical evidence of non-linearity, Section 4 offers psychophysical and cognitive foundation, Section 5 details the non-linear scoring model, Section 6 and 7 cover the practical implications and step-by-step guide on building a non-linear scorecard, followed by Limitations and Future Work (Section 8).
+Analytic Translation Quality Evaluation (TQE) is widely regarded as a high-precision approach to assessing translation quality.
+It is based on systematic annotation of linguistic and technical errors using a predefined taxonomy, most commonly derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], and supported by structured annotator training.
+Once annotations have been produced, a scoring procedure is applied to convert error patterns into a quantitative quality estimate.
+In operational settings, most scoring schemes adopt an implicit linear scaling assumption: the number of permissible errors is expected to increase proportionally with the word count of the evaluated sample.
+In practice, this typically involves normalizing error counts to a reference sample size (e.g., 1,000 or 2,000 words) and comparing them against a user-defined threshold, such as allowing no more than five minor errors per 1,000 words.
+The resulting normalized error density is then used either to make an acceptability decision (pass/fail) or to derive a continuous quality score.
+[^1] Logrus Global Statistics https://logrusglobal.substack.com/archive.
+While such linear normalization can be adequate for sample lengths near the calibration reference, it introduces systematic distortions when applied to substantially shorter or longer texts [gladkoff2021].
+Practitioners and researchers have repeatedly reported mismatches between linear penalty-based scoring and holistic judgments of translation quality.
+For instance, Waddington’s comparative studies [waddington1999, waddington2001] show that additive points-off models alone correlate poorly with holistic evaluations, whereas combined analytic–holistic approaches can improve agreement.
+Similar effects have been documented in related evaluation domains.
+In speech and video quality assessment, international standards observe saturation effects in opinion scores and recommend comparison-based designs and non-linear mappings (typically logistic) between objective indices and Mean Opinion Scores (MOS) [ituP800, streijl2016, vqeglogistic].
+Within translation quality assurance practice, non-linearity is also reflected through severity weighting and the treatment of critical “showstopper” errors, where a single major defect may outweigh many minor ones [mqm2014, saeJ2450].
+Large-scale machine translation evaluations further show that expert, context-aware severity judgments yield rankings that diverge from simple averaging, providing additional evidence for non-additive scoring [freitag2021].
+Related patterns are well established in usability engineering, where a single catastrophic defect can dominate acceptability decisions [nielsen1994].
+Collectively, these findings suggest that perceived translation quality is not well characterized by linear accumulation of penalties.
+Despite these known limitations, linear error scaling remains common in industry practice, partly due to institutional convention and partly due to the tendency to apply linear extrapolation in numerical reasoning.
+Importantly, when evaluators assess translations without explicitly performing normalization, their judgments often deviate from what a linear model would predict.
+[^2] We provide examples of institutional user reports in Appendix A.
+This motivates the research questions addressed in this work:
+RQ1: How does the maximum number of tolerable MQM errors change as a function of evaluation sample length in real-world translation quality assessment settings?
+RQ2: How can this relationship be operationalized as a calibrated non-linear scoring model that remains interpretable, fair across sample lengths, and compatible with existing MQM-based evaluation workflows?
+To answer these questions, we extend our earlier Multi-Range TQE framework [GladkoffEtAl2024] and argue that tolerance for errors increases with sample length, but at a diminishing rate; equivalently, the acceptable error rate per 1,000 words decreases as text length increases.
+We model this behavior using a logarithmic tolerance function and derive a practical calibration procedure that can be integrated into existing MQM workflows with minimal modification.
+Contributions.
+This paper makes the following contributions:
+  • We provide empirical evidence from three large-scale production evaluation settings addressing RQ1, showing that acceptable error counts grow sublinearly with sample size, consistent with a logarithmic dependence.
+  • We link the observed scaling behavior to established psychophysical and cognitive theories, including the Weber–Fechner law and Cognitive Load Theory, thereby motivating a principled non-linear tolerance function.
+  • We propose a calibrated two-parameter non-linear tolerance model (RQ2) that can be fit from two tolerance points via a one-dimensional root-finding procedure, yielding a practical scoring function for variable-length evaluation.
+  • We show that the proposed model can be integrated into existing MQM-based analytic TQE workflows with minimal modification by replacing constant error tolerance with a length-dependent tolerance function, improving interpretability and supporting more consistent evaluation across sample sizes.
+The remainder of this paper is structured as follows.
+Section 2 reviews analytic TQE and MQM-based scoring.
+Section 3 presents empirical evidence for non-linear tolerance growth.
+Section 4 discusses psychophysical and cognitive foundations.
+Section 5 introduces the proposed non-linear scoring model.
+Sections 6 and 7 discuss practical implications and provide guidance for constructing a non-linear scorecard.
+Section 8 concludes with limitations and directions for future work.
 2 Background
 2.1 Analytic TQE and MQM
-Analytic TQE refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications.
-Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories.
-The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014].
-MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling.
-Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the user’s preference.
-Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict.
-Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score.
-For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words  [^3] https://themqm.org/.
+Analytic Translation Quality Evaluation (TQE) is a structured approach to assessing translation output by identifying, categorizing, and recording observable errors with respect to the source text and relevant specifications.
+In contrast to holistic evaluation methods, which produce a single overall impression of quality, analytic TQE requires systematic annotation of discrete error instances, each assigned to a predefined error type and (typically) a severity level.
+The most widely adopted framework for such structured error annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014].
+MQM defines a hierarchical, extensible typology with high-level error dimensions such as Accuracy and Linguistic Conventions (formerly Fluency), and more fine-grained categories including Mistranslation, Omission, Grammar, and Spelling.
+In practice, many industry evaluation schemes implement MQM directly or adopt customized subsets tailored to specific content types, client requirements, or quality objectives.
+Following annotation, a scoring procedure is applied to an evaluation sample—a defined portion of a translation job—to convert the set of annotated errors into either (i) a numeric quality score or (ii) an acceptability decision (pass/fail).
+MQM-based evaluation therefore consists of two complementary components: (i) an error typology specifying which error categories and severity levels are annotated, and (ii) a scoring model that maps these annotations to severity-weighted penalty points and an overall score.
+For comparability across texts of different lengths, scores are typically expressed relative to a fixed reference sample size, most commonly 1,000–2,000 words.
+[^3] https://themqm.org/.
 2.2 Linear Scoring Models and Their Limits
-The dominant approach to scoring in translation quality evaluation is the linear model.
-Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size so that assessments are comparable across jobs of different lengths.
-For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
-Although this proportional scaling appears objective, it fails to reflect how human content users actually perceive quality.
-Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450].
-However, this tolerance wanes as the document length grows further and the errors begin to compound.
-For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance.
-Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
-Building on our earlier work on Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
-  • micro-range (<250<250 words): For very short samples (under 250 words), traditional quality scoring methods are unreliable.
-As shown in [gladkoff2021], quality scores for such brief samples exhibit extremely wide confidence intervals—meaning the range of possible true scores is so large that the measurement is imprecise—making individual measurements meaningless for decision-making.
-    This limitation is particularly important for automatic Quality Estimation (QE), especially with the practice of sentence- or segment-level scoring.
-Recent industry findings confirm that these fine-grained predictions do not align with how humans perceive quality.
-Bane [baneATA2025] highlighted the repeated failures of sentence- or segment-level QE in high-volume translation, where automated scores fail to identify all major, and even critical, errors, making the method unreliable.
-As discussed in [gladkoff2025substack], attempting to derive meaningful scores from very small samples or isolated sentences or segments simply does not work.
-    Instead, for micro-range samples, Statistical Quality Control (SQC) methods offer a more statistically robust and risk-aware framework for interpreting quality signals, explicitly framing the producer’s and consumer’s risk associated with very limited observations.
-  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a user-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
-  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another model is needed: a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
-The present study concentrates on the entire span of statistically reliable measurements in both the meso-range and the macro-range and supplies both psychophysical and empirical evidence for adopting a logarithmic scoring model across these ranges.
-Here “statistically reliable” is meant in the technical sense—i.e., estimates with reasonably narrow confidence intervals —the usual sampling-uncertainty bands around an estimate that shrink as sample size grows.
+The dominant scoring approach in analytic TQE is based on linear length normalization.
+Under this assumption, the severity-weighted penalty total (APT, Absolute Penalty Total; i.e., the sum of penalty points assigned to annotated errors) is scaled proportionally to a reference word count so that results can be compared across evaluation samples of different sizes.
+For example, a tolerance of five minor errors (one point each) per 1,000 words corresponds to 2.5 penalty points in a 500-word sample and 10 penalty points in a 2,000-word sample.
+If major errors receive a $5\times$ severity multiplier, then a single major error consumes the full five-point allowance for a 1,000-word sample.
+Although proportional scaling provides a simple and operationally convenient normalization rule, it does not fully capture how translation quality is perceived in practice.
+Empirical and practitioner-oriented work suggests that evaluators apply stricter expectations to very short texts, where individual errors are salient, while showing greater tolerance for longer passages [waddington1999, waddington2001, mqm2014, saeJ2450].
+At the same time, tolerance does not increase indefinitely with length: as documents grow, error accumulation and compounding effects contribute to declining acceptability.
+For instance, a single major error may be sufficient to fail a one-page sample, yet a seven-page sample would not typically be judged acceptable with seven major errors, even though proportional scaling would imply this threshold.
+Such judgments indicate that the relationship between sample length and acceptable error count is fundamentally non-linear.
+Building on the Multi-Range framework proposed in [GladkoffEtAl2024], we distinguish three regimes of evaluation sample size:
+  • Micro-range (<250<250 words): For very short samples, analytic scoring exhibits high sampling uncertainty.
+As shown in [gladkoff2021], estimates derived from such brief segments yield wide confidence intervals, limiting their interpretability for operational decision-making.
+    This limitation is also relevant for sentence- or segment-level Quality Estimation (QE).
+Recent industry observations suggest that fine-grained QE predictions may not consistently reflect document-level acceptability, particularly when critical errors occur sparsely or depend on broader context [baneATA2025, gladkoff2025substack].
+These findings motivate treating micro-range scoring as a distinct regime in which statistical uncertainty dominates, and where Statistical Quality Control (SQC) methods can provide a more risk-aware interpretation of limited observations.
+  • Meso-range ($\approx 250\!-\!3{,}000$ words): Within this range, scoring based on a reference-length normalization can be operationally effective.
+In particular, a linearly calibrated scoring model anchored to a user-defined tolerance in severity-weighted penalty points (APT) at a reference word count is commonly used.
+  • Macro-range (>3,000>3{,}000 words): For longer samples, human judgments increasingly diverge from linear extrapolation calibrated in the meso-range.
+In this regime, a non-linear tolerance function—in particular, a logarithmic scaling curve—is better suited to model perceived quality across both meso- and macro-range texts without requiring multiple piecewise linear approximations.
+The present study focuses on the meso- and macro-range, where analytic measurements are statistically more stable and practically meaningful.
+Here, “statistically reliable” refers to estimates with sufficiently narrow confidence intervals, i.e., sampling-uncertainty bounds that shrink as the evaluated sample size increases.
 [^4] For intuition, a simple 95% Wald interval for a proportion is $p\pm 1.96\sqrt{p(1-p)/n}$ [gladkoff2021].
-For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or its closely related Agresti–Coull “plus-four” approximation—provides markedly better coverage [AgrestiCoull1998].
+For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or the closely related Agresti–Coull “plus-four” approximation—provides improved coverage [AgrestiCoull1998].
 Wilson 95% CI: $\frac{\hat{p}+\frac{z^{2}}{2n}\pm z\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},$ with z=1.96z\!=\!1.96; the Agresti–Coull “plus-four” uses $\tilde{p}=(x+\tfrac{z^{2}}{2})/(n+z^{2})$ and $\tilde{p}\pm z\sqrt{\tfrac{\tilde{p}(1-\tilde{p})}{n+z^{2}}}$ [AgrestiCoull1998].
-When linear scoring is applied indiscriminately across all ranges, it produces systematic friction.
-Users and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples.
-Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment.
-Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably.
-[^5] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
-The persistent use of linear models—despite long-recognized shortcomings—stems largely from the absence of simple, theoretically grounded alternatives.
-This paper aims to fill that gap by presenting a non-linear scoring model and providing empirical and theoretical support for it, grounded in psychophysics and cognitive-load research.
+Applying linear scoring indiscriminately across all sample sizes often leads to systematic friction in evaluation practice.
+Users and linguists may need to override, reinterpret, or replace scores, or maintain multiple length-specific scoring rules for short and long samples.
+A common workaround is to restrict evaluation to a fixed 1,000-word segment; however, this strategy is operationally restrictive and does not address the underlying mismatch between linear scaling and perceived quality.
+Moreover, because confidence interval width decreases approximately as $1/\sqrt{n}$—with an additional finite-population correction when the evaluated sample constitutes a substantial fraction of the document—variation in sampled word count can materially affect estimation uncertainty.
+[^5] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to the simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
+The persistence of linear scoring practices, despite these limitations, can be attributed in part to the lack of simple, theoretically grounded alternatives that can be deployed within established MQM workflows.
+The present work addresses this gap by proposing a calibrated non-linear scoring model, supported by both empirical evidence and theoretical motivation drawn from psychophysics and cognitive-load research.
 3 Empirical Evidence of Non-Linearity
-To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large institutional users (anonymized).
-Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
+To examine how perceived translation quality varies with evaluation sample length, we conducted an empirical elicitation study with Quality Managers from three large institutional translation programs (anonymized).
+Each organization operates a high-volume multilingual workflow and applies mature, metric-driven TQE procedures to support routine quality monitoring and acceptance decisions.
 3.1 Survey Design
-The extended calibration questionnaire described in [GladkoffEtAl2024] was reused, asking survey respondents to state acceptable error counts for multiple evaluation sample sizes without performing linear extrapolation.
-This method again proved effective in eliciting a latent non-linear response curve.
-A detailed calibration questionnaire was designed to capture Quality Managers’ intuition about acceptable error tolerance—operationalized as the maximum number of severity-weighted penalty points (APT) they would still pass—across a range of evaluation sample sizes.
-Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work.
-Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds.
-To counteract this bias, the institutions’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
-The questionnaire asked the following questions:
+We reused the extended calibration questionnaire introduced in [GladkoffEtAl2024].
+The instrument elicits respondents’ tolerance judgments for multiple sample sizes while explicitly discouraging linear extrapolation, thereby enabling the recovery of latent non-linear tolerance functions.
+The questionnaire was designed to capture Quality Managers’ intuitive notion of error tolerance, operationalized as the maximum number of severity-weighted penalty points (APT) that would still be considered acceptable for a given sample size.
+The key methodological challenge was controlling for respondents’ prior exposure to standard length-normalized scoring schemes, which can lead professionals to default to proportional mental scaling.
+To mitigate such anchoring effects, respondents were instructed not to compute tolerances from existing formulas or reference thresholds, but instead to report what their organization would consider acceptable based on professional judgment and operational experience.
+The questionnaire contained the following prompts:
 Intuitive Error Tolerance Questions Intuitively, how many minor errors would you accept in a 1-page sample?
 Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages?
-For a 3-page text, how many minor errors seem tolerable—just based on your experience?
+For a 3-page text, how many minor errors seem tolerable—based on your experience?
 Do not calculate—just estimate: how many minor errors would you allow in 4 pages?
-Based on gut feeling, how many minor errors would still be okay in a 5-page sample?
+Based on your intuition, how many minor errors would still be acceptable in a 5-page sample?
 Without applying your formal model, what number of minor errors seems acceptable in 10 pages?
-Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
+Without referencing formulas, what minor error count would feel acceptable for 20 pages?
 Elicitation controls.
-To reduce linear anchoring, page sizes were presented in randomized order without showing prior answers; two page sizes were repeated at the end to probe intra‑rater stability.
-Respondents were instructed explicitly not to prorate from known thresholds and to answer from professional judgment alone.
-The questionnaire deliberately focused on minor errors for two key reasons.
-First, higher-severity errors (major and critical) are too infrequent for reliable statistical analysis.
-Second, as many studies note, their impact on quality perception is significantly different and more pronounced.
-This specific focus is practical, as standard industry practice allows the severity of any error to be expressed as an equivalent number of minor errors.
-While institutions used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.
+To reduce anchoring and promote independent judgments across lengths, page sizes were presented in randomized order and prior responses were not displayed.
+Two page sizes were repeated at the end of the questionnaire to probe intra-rater consistency.
+Respondents were again explicitly instructed not to prorate from known thresholds and to rely on professional judgment rather than calculation.
+The elicitation focused on minor errors for two reasons.
+First, higher-severity errors (major and critical) occur relatively infrequently, which makes them less suitable for robust quantitative elicitation at the sample sizes considered.
+Second, their perceptual and operational impact differs qualitatively, as a small number of high-severity defects may dominate acceptability decisions.
+In industry practice, severity-weighted scoring allows the effect of any error to be expressed as an equivalent number of minor errors, making minor-error tolerance a practical proxy for overall tolerance modeling.
+Respondents expressed sample length either in pages or in word counts. For comparability across institutions, all answers were normalized to pages using a conversion factor of 250 words per page.
 3.2 Results
-The most striking finding was a clear non-linear pattern in error tolerance.
-Freed from the habit of linear extrapolation, respondents indicated that while they would accept more total errors in longer texts, their tolerance grew far more slowly than the sample size.
-Instead, the tolerance curve rose steeply at first and then gradually flattened.
-This pattern was consistent across all three anonymized institutions, despite differences in their baseline quality standards (the specific error thresholds and performance targets they use) and content types.
-Figures 13 illustrate the raw responses.
-For each, the horizontal axis represents the sample size in pages, and the vertical axis shows the maximum number of minor errors that the respondent deemed acceptable.
-In one case, both minor and major error series were collected.
-Figure 1: Institution 1 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
-Figure 2: Institution 2 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of minor errors.
-Figure 3: Institution 3 raw questionnaire responses. X‑axis: pages ($250$ words/page); Y‑axis: maximum allowed number of errors (two series shown, minor and major errors).
-None of the institutions’ tolerance curves followed a straight line.
-Although each organization had a different baseline for error tolerance—meaning they accepted different absolute numbers of errors for the same sample length—the overall shape of their tolerance functions was similar and best approximated by a logarithmic curve.
-The respondents themselves were surprised by this result.
-When asked to suspend calculation and rely purely on personal feeling, they quickly identified that their personal perception of what was ‘acceptable’ was non-linear.
-They became aware that their tolerance per page was higher in shorter samples and lower in longer ones, contrary to what linear extrapolation would dictate.
-This phenomenon aligns with everyday experience.
-In very short samples, a single error can be jarring and dominate the user’s entire impression of quality.
-In longer documents, however, the dynamic shifts: while the perceived impact of any one mistake diminishes, the overall experience is instead degraded by the steady accumulation of minor issues, which erodes trust and creates a sense of general sloppiness.
-This degradation impacts more than just a pleasant ”reading experience”; it affects the content’s practical utility, impairing a user’s ability to perform a task, trust instructions, or operate a product correctly.
-Ultimately, these findings confirm that the linear scoring model does not reflect how humans judge translation quality: acceptability is not a linear function of error count, because the perceived impact of an error depends on how much text has already been processed—early mistakes loom larger, while later ones add less per unit but accumulate as the sample grows.
-Consistent with Fig.
-7 in our pilot study from [GladkoffEtAl2024], our findings show that linear models overestimate acceptable error counts in the macro-range by up to 50%, while a logarithmic model closely tracks expert judgment throughout the 1-20 page span.
-Based on this consistent pattern, this paper hypothesizes that human tolerance to translation errors grows logarithmically with the sample size.
-The rest of this paper explores the theoretical underpinnings of this phenomenon and proposes a formal scoring model that better aligns with real-world perceptions.
-From Figures 13, it is already clear that the data points do not fall on a straight line.
-The logarithmic curve, in contrast, follows the shape of the expert judgments closely.
-For readers who want the formal evidence behind Fig.
-3: we detail the linear vs logarithmic fits in Appendix B.
+Across all three institutions, responses exhibited a consistent non-linear tolerance pattern.
+Respondents indicated that longer samples permit a higher absolute number of errors, but that this permissible count grows substantially more slowly than sample length.
+Empirically, tolerance increased steeply for short samples and gradually saturated for larger samples.
+This qualitative shape was observed despite differences in each institution’s baseline standards, performance targets, and content types.
+Figures 13 report the raw elicitation results.
+The horizontal axis represents sample length in pages and the vertical axis represents the maximum number of errors considered acceptable.
+In one institution, tolerance judgments were collected for both minor and major error series.
+Figure 1: Institution 1 raw questionnaire responses. X-axis: sample size in pages ($250$ words/page); Y-axis: maximum acceptable number of minor errors.
+Figure 2: Institution 2 raw questionnaire responses.
+X-axis: sample size in pages ($250$ words/page).
+Y-axis: fail thresholds (maximum acceptable counts) for four elicited series: minor and major errors, each for TERM/ACC and STYLE categories.
+Figure 3: Institution 3 raw questionnaire responses. X-axis: pages ($250$ words/page); Y-axis: maximum allowed number of errors (two series shown, minor and major errors).
+None of the observed tolerance functions were well described by a straight-line relationship.
+Although institutions differed in absolute tolerance level, the response curves shared a similar concave shape and were well approximated by a logarithmic form.
+Notably, respondents reported that the elicitation process itself made the non-linearity salient: when asked to suspend explicit calculation, they recognized that their implicit acceptance criteria assign disproportionately high weight to errors in short samples, while additional errors in longer texts contribute less marginal impact.
+These results suggest that acceptance judgments depend not only on the absolute number of errors, but also on how errors accumulate across increasing text length.
+As a consequence, tolerance does not scale proportionally with sample size, resulting in non-linear growth in acceptable error counts.
+Consistent with the pilot results reported in [GladkoffEtAl2024], the present findings indicate that linear scaling substantially overestimates tolerable error counts in the macro-range, in some cases by up to 50%, whereas a logarithmic tolerance function tracks expert judgments more closely over the 1–20 page range.
+Based on this convergent pattern, we hypothesize that human tolerance for translation errors grows approximately logarithmically with sample size.
+The remainder of this paper examines theoretical explanations for this relationship and introduces a calibrated scoring model that operationalizes it within MQM-based evaluation practice.
+From Figures 13, it is evident that the elicited tolerance points do not follow a linear trend.
+For readers interested in formal model comparison, Appendix B provides quantitative details on linear versus logarithmic fits, including the results underlying the trend observed in Figure 3.
+We note that these results reflect expert elicitation rather than controlled behavioral experiments; nevertheless, the consistent concave trend across institutions provides strong evidence of non-linear tolerance scaling in operational settings.
 4 Psychophysical and Cognitive Foundations
-Empirical findings from these anonymized institution interviews show that perceived translation quality is not a simple linear function of error count or text length.
-Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950].
-This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
+The elicitation results in Section 3 suggest that perceived translation quality does not vary linearly with either sample length or the total number of annotated errors.
+This section provides theoretical motivation for such non-linearity by drawing on two complementary perspectives: classical psychophysical scaling and Cognitive Load Theory.
+We additionally relate these accounts to document-level MT evaluation findings, where context and aggregation are known to influence human judgments and agreement patterns [castilho-2020-page, grundkiewicz-etal-2021-user, vernikos-etal-2022-embarrassingly, kim-2025-context].
+Scope and interpretive caveat.
+Translation errors are not physical stimuli in the strict psychophysical sense, and we do not claim that quality judgments obey a literal Weber–Fechner law.
+Rather, psychophysics and cognitive theories offer principled functional explanations for why human sensitivity to additional defects is typically compressive and why aggregation over longer texts can yield length-dependent acceptability thresholds.
+These perspectives provide a theoretically grounded motivation for replacing proportional (linear) extrapolation with a calibrated sublinear tolerance function in MQM-style analytic scoring.
 4.1 Weber–Fechner Law and Logarithmic Perception
-Psychophysics has long recognized that sensation grows non-linearly with stimulus intensity.
-In his 1834 monograph on weight discrimination, Weber reported that just-noticeable difference (JND) is a constant fraction of the baseline stimulus: “The smallest perceptible increase of weight bears a fixed ratio to the weight already lifted[weber1834, p.
-152].
-Three decades later, Fechner integrated Weber’s proportionality into a mathematical model of subjective sensation, arguing that “if equal stimulus ratios yield equal sensation increments, then sensation must be the logarithm of stimulus magnitude[fechner1860, vol.
-1, p.
-55].
-This logarithmic relation—now known as the Weber–Fechner law—was the first quantitative statement of a compressive perceptual scale; an authoritative English translation is provided in [fechner1966].
-Intuition.
-Applied to translation quality evaluation, the law would predict that the subjective penalty of an additional error is highest when the text is nearly error-free and decreases as errors accumulate.
-In other words, the first mistranslation in a short paragraph feels much more jarring than the fifteenth typo in a 20-page report.
-Subsequent psychophysical research refined, but did not overturn, this insight.
-Stevens showed that many modalities follow a power function $S=k\,I^{n}$ with exponents $n<1$ for brightness, loudness, and heaviness [stevens1957].
-Although $n=0$ is not itself logarithmic, the power-law family $I^{n}$ is asymptotically log-like:
+Psychophysics has long established that subjective sensation typically grows non-linearly with stimulus intensity.
+Weber observed that the just-noticeable difference (JND) is approximately proportional to the baseline stimulus magnitude [weber1834].
+Fechner subsequently formalized a quantitative model in which equal stimulus ratios correspond to equal increments in sensation, implying a logarithmic relationship between physical intensity and perceived magnitude [fechner1860].
+This principle, now commonly referred to as the Weber–Fechner law, is a canonical example of compressive perceptual scaling (see also [boring1950, fechner1966]).
+Implication for TQE.
+As an analogy for translation quality evaluation, compressive scaling predicts diminishing marginal perceptual impact of additional errors.
+That is, the subjective penalty associated with one more defect is expected to be larger when the text is nearly error-free and smaller when multiple errors have already been encountered.
+This prediction is consistent with the empirical pattern observed in Section 3, where tolerance rises rapidly for short samples and flattens as sample length increases.
+Subsequent psychophysical work refined this relationship while preserving its central insight.
+Stevens proposed that many perceptual modalities follow a power function $S=kI^{n}$, often with exponents $n<1$, yielding concave response curves consistent with sublinear growth [stevens1957].
+Moreover, the logarithmic function can be viewed as a limiting case of a power law:
 $\lim_{n\to 0}\frac{I^{\,n}-1}{n}\;=\;\ln I.$
-Over finite dynamic ranges, exponents in the $0.2\text{–}0.4$ band produce curves that closely approximate a logarithmic profile.
-Over typical stimulus ranges, power exponents $n\in[0.2,0.4]$ are empirically indistinguishable from a log on plots like ours, so a logarithmic tolerance is the parsimonious choice here.
-Comprehensive reviews such as [gescheider1997] and the signal detection synthesis of [green1966] confirm that compressive (sublinear) mapping is the norm in vision, audition and touch.
-Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales.
-Treating translation errors as discrete ‘stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
-4.2 Cognitive Load and Accumulated Impact of Errors
-Whereas psychophysics explains the decrease in sensitivity to additional stimuli, cognitive psychology accounts for how these stimuli collectively strain mental resources.
-The theory of limited working memory, formalized as Cognitive Load Theory (CLT), provides a parallel foundation for non-linear perception of translation quality.
-CLT was introduced by Sweller in the late 1980s as a way to explain why instructional design must respect the cognitive limitations of learners [sweller1988].
-It builds on earlier findings such as Miller’s famous observation that the number of discrete items a person can hold in short-term memory is limited to approximately seven, plus or minus two [miller1956].
-As Sweller notes, “Our capacity to process novel information is extremely limited, and when this limit is exceeded, learning and comprehension suffer[sweller1988, p.
-257].
-In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources.
-One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity.
-This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality.
-Importantly, this degradation does not occur linearly.
-Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency.
-Studies in instructional design and reading comprehension (e.g.
-[mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
-Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001].
-This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content.
-Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework.
-Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption.
-Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
-4.3 Perception and Cognition Converge on Logarithmic Scoring
-In summary, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically.
-A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors.
-This synthesis underpins the non-linear scoring model proposed in the next section.
+Over practical dynamic ranges, small-exponent power functions can approximate logarithmic behavior closely.
+A logarithmic tolerance curve therefore provides a parsimonious functional form for modeling diminishing marginal sensitivity in quality judgments.
+Comprehensive psychophysical treatments and signal detection theory syntheses likewise emphasize compressive mappings as common across perceptual domains [gescheider1997, green1966].
+Taken together, the psychophysical literature motivates the hypothesis that translation errors, treated as discrete “quality-relevant events,” are integrated by human evaluators using a compressive internal scale.
+This provides one theoretical justification for sublinear growth in acceptable error counts as evaluation samples become longer.
+4.2 Cognitive Load and Cumulative Disruption
+Whereas psychophysical scaling accounts for diminishing marginal sensitivity to repeated stimuli, cognitive theories emphasize how repeated disruptions impose increasing demands on limited processing resources.
+Cognitive Load Theory (CLT) formalizes the idea that working memory capacity is constrained and that comprehension degrades when cognitive demands exceed these limits [sweller1988], consistent with foundational observations on short-term memory capacity [miller1956].
+In translation evaluation, each error can function as a localized disruption that requires attention, reinterpretation, or correction by the reader.
+While isolated minor errors may be tolerated, repeated disruptions can accumulate and reduce processing fluency, increasing perceived effort and lowering confidence in the text.
+As a result, acceptability can decline non-linearly with total error accumulation even when errors are individually low-severity.
+Related effects are documented in document-level MT evaluation: presenting broader context can change judgments and alter agreement patterns relative to sentence-level setups [castilho-2020-page, grundkiewicz-etal-2021-user], and document-level aggregation can capture quality factors that are not reliably reflected by sentence-level scoring alone [vernikos-etal-2022-embarrassingly].
+Recent analyses further suggest that “context” effects may be pervasive yet heterogeneous, complicating attempts to infer document-level acceptability from local judgments [kim-2025-context].
+CLT also provides a plausible interpretation of length effects reported in applied translation assessment: as the expected cognitive demands of processing a long document increase, evaluators may adopt stricter acceptance thresholds to reduce the risk of cumulative disruption, consistent with divergences between analytic scoring and holistic judgments in longer samples [waddington1999, waddington2001].
+4.3 Implications for Length-Dependent Tolerance Modeling
+In summary, psychophysical scaling and cognitive-load considerations converge on a shared expectation: tolerance should increase with sample length, but at a diminishing rate.
+A length-dependent logarithmic tolerance function therefore offers a compact way to reconcile (i) diminishing marginal perceptual impact of individual errors with (ii) increasing cumulative processing costs as disruptions accumulate across longer texts.
+This theoretical synthesis motivates the non-linear scoring model introduced in Section 5.
 5 Proposed Non-Linear Scoring Model
-This paper adopts the terminology introduced in or primary study [GladkoffEtAl2024], distinguishing between three classes of scoring models.
-Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”).
-This uses the sample’s penalty rate and is not a calibrated/anchored score.
-Calibrated Linear: This scoring model maps a user/institution’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
-Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve.
-Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
-Given both the empirical evidence and the theoretical justifications presented, this paper proposes a more nuanced application of scoring models rather than a complete replacement of the linear approach.
-While the traditional linear model is often considered sufficient for the meso-range (texts up to 3,000 words), our findings indicate this is precisely where its limitations can be most acute.
-For example, as shown later in this paper, a linear model calibrated for a 1,000-word reference sample loses significant accuracy (over 20%) when applied to a 2,000-word sample.
-Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4).
-Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated.
-666‘Calibration’ for the raw score simply means choosing the user’s pass/fail cut-point on the raw penalty-rate scale.
-Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words).
-A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$.
-No calibration mapping to a 0–100 quality scale is involved.
-Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring.
-To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
-This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages.
-By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture.
-Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system.
-5.1 Review the Limitations of Linear Extrapolation
-Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size.
-For example, if a user accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc.
-However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
-As noted above, surveyed institutional users describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound.
-In practice, however, evaluators do not extend that allowance proportionally.
-When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors.
-In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length.
-What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole.
-This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
+We adopt the terminology introduced in our primary study [GladkoffEtAl2024], which distinguishes three classes of scoring approaches for MQM-style analytic Translation Quality Evaluation (TQE):
+Raw score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC},$ where $\alpha$ is a unit scaling factor (e.g., $\alpha=1000$ for “per 1,000 words”).
+This formulation reflects the sample’s observed penalty rate but is not anchored to an institutional tolerance threshold and does not yield a calibrated quality scale.
+Calibrated linear model: a scoring model that maps an institution’s acceptability tolerance (pass/fail threshold) onto a normalized 0–100 scale, enabling interpretability and comparability within a fixed reference range.
+Calibrated non-linear model: a scoring model that replaces proportional scaling with a length-dependent tolerance function, modeled here as a logarithmic curve.
+Relative to a linear rule anchored at a reference size, the non-linear model is more permissive for short samples and more stringent for long samples; total tolerance still increases with length, but at a diminishing rate.
+Given the empirical results in Section 3 and the theoretical motivation in Section 4, we do not argue for eliminating linear scoring entirely.
+Rather, we propose a principled regime of applicability: linear scoring can remain useful near its reference calibration size, while a non-linear tolerance function is required for robust length generalization.
+In particular, even within the meso-range (up to approximately 3,000 words), linear extrapolation can introduce substantial distortion when applied outside its reference vicinity.
+For example, as shown in Section 5.4, a linear model anchored at a 1,000-word reference point can exceed a 20% relative deviation when applied to a 2,000-word sample.
+Accordingly, we recommend using the logarithmic model whenever evaluation sample sizes are expected to vary by more than approximately $20\%$ from the reference word count, i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4.
+Raw scoring and calibrated linear scoring remain viable within restricted settings, provided that evaluation samples consistently remain near the calibration length.
+[^6] For raw scoring, “calibration” amounts to selecting a pass/fail cutoff on the penalty-rate scale.
+Let $r_{\text{thr}}$ denote the maximum acceptable penalty rate (e.g., 5 points per 1,000 words).
+A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes iff $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$.
+No mapping to a normalized 0–100 scale is involved.
+Some implementations rescale the penalty rate into a 0–100 “raw score” via $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this inherits the same limitations of proportional length scaling.
+To obtain scores that remain comparable across varying evaluation sizes, a calibrated model is required (linear near the reference size, non-linear over wider ranges).
+This section introduces the mathematical form of the proposed logarithmic tolerance model, describes its anchoring and calibration procedures, and characterizes the regime in which linear approximation remains sufficiently accurate.
+By extending MQM scoring with a length-dependent tolerance function, the proposed approach enables analytic evaluation to better align with document-level expert judgments under variable-length sampling conditions.
+5.1 Limitations of Linear Extrapolation
+Linear scoring models assume that acceptable error counts increase proportionally with sample length.
+For example, if an institution accepts 5 minor errors in a 1,000-word sample, linear scaling permits 10 minor errors in 2,000 words and 15 in 3,000 words.
+However, as shown in Sections 3 and 4, this assumption is not consistent with elicited expert tolerances across variable-length samples.
+In practice, institutional guidance may appear linear when expressed informally (e.g., “approximately one serious error per page”), yet observed acceptance judgments are typically sublinear.
+When sample size increases to multiple pages, evaluators do not generally accept a strictly proportional increase in severe errors; instead, tolerance per unit length decreases with text length.
+This indicates that proportional extrapolation can systematically overestimate acceptable error totals for longer samples and underestimate the impact of errors in shorter samples.
 5.2 Proposed Mathematical Model of Acceptable Errors
-This paper proposes the following logarithmic function to model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$, expressed in words:
+We model the maximum acceptable number of errors $E(x)$ as a function of sample size $x$ (measured in words) using the following logarithmic tolerance function:
 $E(x)\;=\;a\,\ln(1+b\,x),\qquad a>0,\;b>0.$ (1)
 Throughout, $\ln$ denotes the natural logarithm.
-If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged.
-This unit invariance ensures figures shown in pages and formulas written in words are consistent.
-The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens.
-This curve shape effectively combines both theoretical expectations and observed data.
-To clarify the functional form, we adopt the intercept-free specification $E(x)=a\,\ln(1+bx)$.
-This choice enforces the natural boundary condition $E(0)=0$ and provides explicit control over curvature through the parameter $b$.
-This formulation is distinct from the “logarithmic trendline” $c+k\ln x$ commonly implemented in spreadsheet software, which introduces a free intercept term and does not pass through the origin.
-Consequently, the latter cannot be anchored to calibration points in an equivalent manner.
-[^7] A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope].
-LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution.
-In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size.
-The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
-Other scoring models may be devised for specific applications.
-Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here.
-Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear.
-For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations.
-Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$.
-Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450].
-(One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper.
-They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely.
+The function satisfies $E(0)=0$, increases rapidly for short samples, and flattens gradually as sample length increases, matching the concave tolerance trend observed empirically.
+If sample size is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $Wb$, while $a$ remains unchanged.
+This unit invariance ensures consistency between page-based visualizations and word-based formulations.
+We adopt the intercept-free specification $E(x)=a\ln(1+bx)$ to enforce the natural boundary condition $E(0)=0$ and to provide explicit control of curvature via parameter $b$.
+This differs from the “logarithmic trendline” form $c+k\ln x$ commonly used in spreadsheet software, which introduces a free intercept term and does not pass through the origin.
+As a result, the spreadsheet form is not directly anchorable to tolerance points in the same manner.
+Other scoring models may be developed for specific applications.
+Importantly, however, severity weighting within a fixed sample is distinct from the length-dependent non-linearity considered here.
+Even an exponential ladder of severity multipliers affects how errors are aggregated within a sample, but it does not, by itself, induce a non-linear relationship between allowable error totals and sample size.
+For example, ATA certification scoring is an analytic, points-off scheme tailored to short, fixed-length exam translations, with steep severity scaling but no tolerance curve $E(x)$ varying with length.
+Similarly, SAE J2450 assigns fixed category weights and sums penalties additively within a sample [saeJ2450].
+Such frameworks introduce non-uniformity across severities, but not the sample-size dependent tolerance function required to keep pass/fail criteria comparable across widely varying evaluation lengths.
 5.3 Anchoring and Calibration Procedures
-This paper distinguishes between two steps in adapting a scoring model to operational use:
-  • Anchoring defines the maximum acceptable error limit for a standard amount of text.
-This tolerance is expressed in penalty points based on a user’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
-  • Calibration takes the absolute penalty point total and maps it onto a normalized, more easily interpretable 0-100 quality scale for use in dashboards or QA reports.
-Feasibility (two‑point).
-Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
+For operational deployment, we distinguish two adaptation steps:
+  • Anchoring: specifying an acceptable error limit at a reference sample size, expressed in severity-weighted penalty points (e.g., “no more than 5 points per 1,000 words”).
+  • Calibration: mapping the observed penalty total for a sample into a normalized 0–100 quality scale, supporting interpretability in dashboards and reporting.
+Two-point feasibility. Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
 $r:=\frac{E_{1}}{E_{0}},\qquad\rho:=\frac{x_{1}}{x_{0}},\qquad g(b):=\frac{\ln(1+bx_{1})}{\ln(1+bx_{0})}.$
@@
 $\boxed{\ \min\{1,\rho\}\;<\;r\;<\;\max\{1,\rho\}\ }.$
-Intuitively, the longer sample must permit more errors but sublinearly; the shorter must permit fewer errors but not in direct proportion to length.
-The model can be easily calibrated using two known tolerance points. Suppose we have:
-  • The primary anchor point: This is the main tolerance defined during the Anchoring step, such as a maximum of 5 errors ($E_{0}=5$) for a reference sample size of 1,000 words ($x_{0}=1000$).
-  • A second, empirical data point: This point is derived from observing the user’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
-Two‑point calibration (numerical).
+Intuitively, the longer sample must permit more errors, but sublinearly, relative to its length increase.
+Two-point calibration (numerical).
 Let $r=E_{1}/E_{0}$ and define $f(b)=\ln(1+bx_{1})-r\,\ln(1+bx_{0})$.
-Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$ by the intermediate value theorem.
-Near $b=0$, $f(b)\approx b(x_{1}-rx_{0})$, so the sign at the origin is $\mathrm{sgn}(x_{1}-rx_{0})$.
-As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$; if $r>1$ the limit is -$\infty$, and if $r<1$ it is $\infty$, guaranteeing a sign change and thus a unique root.
-Solve $f(b)=0$ with a robust one‑dimensional method (e.g., bisection/Brent); then set
+Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$.
+The parameter $b^{\star}$ can be obtained with a robust one-dimensional root-finding method (e.g., bisection or Brent), after which
 $a^{\star}=\frac{E_{0}}{\ln(1+b^{\star}x_{0})}.$
-See  Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
+See Calibration from Two Points (Numerical) for a one-dimensional root-finding implementation.
 Illustrative values.
-Feasibility.
-Ensure $r=E_{1}/E_{0}$ lies strictly between $1$ and $\rho=x_{1}/x_{0}$, i.e., $\min\{1,\rho\}<r<\max\{1,\rho\}$, before solving for $b$.
-Solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving the calibrated curve
+For the example anchor points $(x_{0},E_{0})=(1000,5)$ and $(x_{1},E_{1})=(250,2)$, solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving
 $E(x)\;=\;3.688\,\ln\!\bigl(1+0.00288\,x\bigr).$
-Fallback for multiple points.
-If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available—or if the two chosen points violate the feasibility condition—estimate $(a,b)$ by constrained least squares:
+Fallback for multiple points. If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available, estimate $(a,b)$ via constrained least squares:
 $\min_{a>0,\;b>0}\ \sum_{i=1}^{n}\Bigl[E_{i}-a\,\ln\!\bigl(1+b\,x_{i}\bigr)\Bigr]^{2}.$
-Given any $b_{0}>0$, a fast 1‑D search over $b$ with the closed‑form $a(b)=\dfrac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$ works well (then refine both jointly if desired).
-This approach allows the model to be tailored to specific content types, organizational standards, or language pairs while still operating under a common mathematical framework.
-For instance, the error tolerance can be precisely defined for each case, ranging from extremely strict for software UI text, to moderately strict for marketing copy and documentation, and significantly more lenient for internal communications or user-generated content.
-For small $b$, $\ln(1+b\,x)\approx b\,x$, so the model reduces locally to a linear tolerance rule; this explains why a calibrated linear slope performs acceptably only in the vicinity of its reference sample size.
-If $x$ is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $W\cdot b$; $a$ is unchanged.
+Given any $b_{0}>0$, a fast one-dimensional search over $b$ with closed-form
+$a(b)=\frac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$
+provides an efficient initialization (followed by optional joint refinement of $(a,b)$).
+For small $b$, $\ln(1+bx)\approx bx$, implying that the model is locally linear near the reference size.
+This explains why calibrated linear scoring can be adequate only in a neighborhood around its anchor point.
 Figure 4: Three interpolations.
-The spreadsheet curve $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
-5.4 Near-Reference Fidelity Interval for Linear Approximation (±20%)
-In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used for convenience:
+The spreadsheet trendline $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
+5.4 Near-Reference Fidelity Interval for Linear Approximation ($\pm 20\%$)
+In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used:
 $E_{\mathrm{lin}}(x)\;=\;E_{\log}(x_{\mathrm{ref}})\,\frac{x}{x_{\mathrm{ref}}},\qquad E_{\log}(x)=a\,\ln\!\bigl(1+bx\bigr).$
-To quantify where this linear approximation is acceptable, we define the ±20% fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
+To characterize where this linear approximation remains acceptable, we define the $\pm 20\%$ fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
 $\left|\frac{E_{\mathrm{lin}}(x)}{E_{\log}(x)}-1\right|\leq 0.20.$
-Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the two boundary points admit a closed form via the Derivation (sketch).
-Set $E_{\rm lin}(x)=(x/x_{\rm ref})\,E_{\rm log}(x_{\rm ref})$ and require $E_{\rm lin}(x)=(1\pm\varepsilon)E_{\rm log}(x)$.
-Writing $\alpha=\ln(1+bx_{\rm ref})/(bx_{\rm ref})$ and $y=1+bx$, we obtain $\alpha\ln y=(1\pm\varepsilon)(y-1)/y$.
-Rearranging gives $y\,e^{-\frac{1\pm\varepsilon}{\alpha}\,y}=e^{-\frac{1\pm\varepsilon}{\alpha}}$, so
-$y=-\frac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\Bigl(-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\Bigr),$
-which yields the closed form in the next equation. The branch $W_{-1}$ selects the solution near $x_{\rm ref}$.
-Lambert $W$ function:
+Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the boundary points admit a closed form using the Lambert $W$ function:
 $x_{\pm}\;=\;\frac{-\,\dfrac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\!\left(\,-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\right)-1}{\,b\,}$,
-where $x_{-}$ (with the “-” sign) gives the lower boundary (linear is $0.8\times$ the log curve) and $x_{+}$ (with the “++” sign) gives the upper boundary (linear is $1.2\times$ the log curve).
-[^8] The branch $W_{-1}$ is selected to obtain the solution near $x_{\mathrm{ref}}$.
-Let’s look at the numerical intervals for the example calibration. Using the illustrative curve from §5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the ±20% fidelity intervals are:
-  • $\boldsymbol{x_{\mathrm{ref}}=1000}$ words: $x\in[579,\;1460]$ words, width $\approx 881$ words (about $42\%$ to $46\%$ relative to $x_{\mathrm{ref}}$).
-  • $\boldsymbol{x_{\mathrm{ref}}=2000}$ words: $x\in[1307,\;2747]$ words, width $\approx 1440$ words (about $34.7\%$ to $37.4\%$ relative to $x_{\mathrm{ref}}$).
-These calculations illustrate how far one can move away from the reference sample size while still retaining acceptable measurement accuracy.
-The window is relatively narrow: if the reference is 1,000 words, then evaluating a 2,000-word sample already falls outside the ±20% fidelity band and will not be accurate; likewise, if the reference is 2,000 words, applying the same linear rule to a 1,000-word sample is equally unreliable.
-In absolute terms, the ±20% window widens as the reference size increases (from about 0.88 k to about 1.44 k words), but in relative terms it becomes proportionally narrower.
-This quantifies the earlier point: a calibrated linear rule is only reliable near its anchoring point.
-Further away, it becomes too harsh for short samples and too lenient for long ones.
-When jobs regularly fall outside these fidelity intervals, the logarithmic model should be used directly, or the linear rule should be re-anchored to a more suitable reference size.
-5.5 Visual comparison across three sample-size ranges
-Figure 5: Linear model (blue) vs.
-human perception (PFT, red).
-Yellow bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule over‑ or under‑estimates true tolerance.
-Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes.
-In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation word-count).
-The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes.
-This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
-Figure 6: Applicability across evaluation sample sizes.
-SQC (micro, $\lesssim\!250$ words), calibrated linear MQM (meso; reliable only within $\sim\!\pm 20\%$ of anchor), and non-linear MQM (macro).
-Vertical mark at $\sim\!250$ words flags the micro‑range where deterministic curves are unreliable.
-Y‑axis: error tolerance (penalty points).
-In summary, building on the Multi-Range Framework [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics.
-In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
-For the meso- and macro-ranges, the choice of an appropriate scoring model depends not on the absolute length of any single text, but on the expected range of fluctuation in evaluation sample sizes.
-This paper’s findings point to a crucial rule of thumb: if evaluation sample sizes are expected to vary by more than approximately 20% from the anchor point, the non-linear model is required for accuracy.
-A calibrated linear model is only a viable tool when evaluation sample sizes remain consistently within this tight  20% band around the anchor.
-Counter-intuitively, because the logarithmic curve flattens at large values, a linear model can be more applicable in the macro-range than the meso-range, provided its anchor point is also in that higher range.
-Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy.
-This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges.
-Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
+where $x_{-}$ corresponds to the lower boundary and $x_{+}$ to the upper boundary. [^7] The branch $W_{-1}$ yields the solution near $x_{\mathrm{ref}}$.
+Using the illustrative calibration in Section 5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the resulting fidelity intervals are:
+  • $\boldsymbol{x_{\mathrm{ref}}=1000}$: $x\in[579,\;1460]$ words.
+  • $\boldsymbol{x_{\mathrm{ref}}=2000}$: $x\in[1307,\;2747]$ words.
+These intervals quantify the limited range over which a single anchored linear slope remains within 20% relative deviation.
+Outside this range, proportional extrapolation becomes increasingly inaccurate, motivating use of the logarithmic model or re-anchoring the linear model to a more appropriate reference size.
+5.5 Visual comparison across sample-size regimes
+Figure 5: Linear model vs.
+non-linear tolerance.
+The shaded bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule under- or over-estimates tolerance.
+Figure 5 illustrates that once the slope is fixed by a single anchor point, a linear rule diverges substantially from the concave tolerance pattern implied by human judgments at both smaller and larger evaluation sizes.
+In contrast, the logarithmic model remains consistent with the length-dependent tolerance trend across the range of sample sizes considered.
+Figure 6: Applicability across evaluation sizes.
+SQC (micro, $\lesssim 250$ words), calibrated linear scoring (meso; reliable only near anchor), and non-linear scoring (macro).
+Y-axis: tolerance in penalty points.
+Finally, Table 1 summarizes operational guidance for selecting scoring models under varying sampling conditions.
 Scenario Recommended model
-EWC varies within $\pm 20\%$ around the anchor Calibrated linear model (re-anchor if the range shifts)
+EWC varies within $\pm 20\%$ of the anchor Calibrated linear model (re-anchor if the range shifts)
 EWC routinely outside the $\pm 20\%$ band Use the non-linear logarithmic model directly
-Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC); deterministic curves unreliable
-Table 1: Operational guidance: which scoring model to use in practice.
+Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC)
+Table 1: Operational guidance for model selection.
 6 Practical Implications
-To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry.
-Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception.
-Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations.
-Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (user-specific PT/MSV mapping).
+If error tolerance scales non-linearly with sample length—as suggested by the elicitation results in Section 3 and the theoretical considerations in Section 4—then adopting a logarithmic tolerance function has practical implications for translation quality evaluation procedures, quality management policies, and supporting software tooling.
+In particular, replacing proportional length normalization with a calibrated, length-dependent tolerance model can reduce systematic length-induced bias and improve the consistency of acceptability decisions across samples of varying size.
+A length-dependent scoring model may also be relevant for emerging AI-assisted document-level evaluation workflows.
+Contemporary systems can provide holistic judgments, but their outputs are often qualitative and may vary in completeness or granularity across documents.
+A calibrated non-linear scoring model offers a quantitative reference function that could be used to normalize or contextualize such judgments against a stable, institution-specific 0–100 scoring scale (e.g., via a user-specific PT/MSV mapping), thereby facilitating comparability across evaluation settings.
 6.1 Operational Guidelines for Quality Managers
-One immediate benefit of the non-linear model is its ability to inform and standardize quality expectations across variable evaluation sample sizes.
-Quality managers can use the calibrated log function to define clear error tolerance thresholds for different content types or use cases, reducing ambiguity when evaluating translation outputs that differ in word count.
-For example, a quality manager might specify: “Our reference wordcount is 2000 words with a tolerance of 7 minor errors.
-For evaluation sample sizes between 1500 and 3000 words, the standard linear approximation is acceptable.
-Outside that range, the logarithmic model must be used to adjust expectations accordingly.”
-This shift allows quality managers and stakeholders to move beyond one-size-fits-all scoring rules and instead adopt quality scores that reflect actual cognitive effort and perceived severity.
-The result is more fair and interpretable evaluation outcomes, especially when comparing translation outputs across different content types or word counts.
-7 Building your non-linear scorecard
-We list the following steps in this section including model and least-squares fit, from tolerance to a calibrated score, and scorecard mapping, followed by implications for CAT tools and linguistic quality assessment (LQA) automation, improvement of a unified calibrated quality scale, applications beyond human translation, and document-level evaluation.
-[^9] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
-http://www.astm.org/workitem-wk46396
+A direct operational benefit of the proposed model is support for consistent tolerance setting across variable evaluation sample sizes.
+Quality managers can use the calibrated logarithmic function $E(x)$ to define explicit acceptability thresholds as a function of evaluation word count, and to adapt tolerances across content types or use cases without relying on ad hoc proportional extrapolation.
+For example, a quality policy can be expressed as follows: “We anchor tolerance at 2,000 words with a maximum of 7 minor-error equivalents.
+Linear approximation is acceptable within the fidelity interval (approximately 1,500–3,000 words in this configuration).
+Outside this range, tolerance thresholds should be computed using the logarithmic model.”
+This approach enables stakeholders to move beyond a single fixed-length scorecard and instead apply acceptance thresholds that reflect length-dependent judgment behavior.
+The resulting evaluation outcomes are more interpretable and more equitable across texts of different lengths, particularly when quality monitoring requires comparison across heterogeneous content types and evaluation sample sizes.
+6.2 Tooling Implications
+From a tooling perspective, the proposed model is straightforward to integrate into existing MQM-based evaluation pipelines.
+Most production systems already compute the severity-weighted penalty total (APT) and the evaluated word count (EWC); therefore, incorporating non-linear length dependence requires only replacing a constant tolerance threshold with a length-dependent tolerance function $E(x)$.
+Concretely, the acceptability condition becomes
+$\mathrm{APT}\leq E(\mathrm{EWC}),$
+rather than $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$ under proportional scaling.
+Because the core annotation workflow, error typology, and penalty aggregation remain unchanged, the model can be deployed without modifying MQM schemas, annotation tools, or reviewer training protocols.
+The primary implementation changes occur at the scoring layer and in reporting interfaces (e.g., dashboards and QA summaries), which must display tolerances and pass/fail thresholds that vary with sample length.
+In addition, calibration parameters $(a,b)$ can be stored per institution, content type, or language pair, enabling consistent application within a unified scoring framework while preserving user-specific acceptance standards.
+Implementation sketch.
+Algorithm 1 summarizes a minimal integration of length-dependent tolerance into a typical MQM scoring pipeline.
+Algorithm 1 Length-dependent tolerance scoring (non-linear MQM)
+1:Annotated errors with severities; evaluated word count $\mathrm{EWC}$; calibrated parameters $(a,b)$; score mapping function $\mathrm{Score}(\cdot)$
+2:Pass/Fail decision; normalized 0–100 score
+3:Compute severity-weighted penalty total $\mathrm{APT}$ from annotations
+4:Compute tolerance threshold $T\leftarrow a\ln(1+b\,\mathrm{EWC})$
+[…]
+[…]
+7:Compute deviation ratio $d\leftarrow\mathrm{APT}/T$
+8:Compute normalized score $s\leftarrow\mathrm{Score}(d)$
+9:return $\mathrm{Pass},\,s$
+7 Building a Non-Linear Scorecard
+This section provides a practical procedure for constructing and deploying a non-linear MQM scorecard, covering (i) model fitting from tolerance points, (ii) conversion from tolerance to a calibrated score on a 0–100 scale, and (iii) scorecard field computation for operational reporting.
+We further discuss implications for CAT tools and LQA automation, the benefits of a unified calibrated quality scale, applications to AI-generated content, and document-level evaluation considerations.
+[^8] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes MQM scoring models and scoring mechanics in detail.
+http://www.astm.org/workitem-wk46396.
 7.1 Model and least-squares fit
-We model the pass/fail tolerance as a function of sample size $x$ (words) by
+We model pass/fail tolerance as a function of evaluation sample size $x$ (in words) using the logarithmic form
 $E(x)\;=\;a\,\ln\!\bigl(1+b\,x\bigr),\qquad a>0,\;b>0.$ (2)
-Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, estimate $(a,b)$ by least squares
+Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, the parameters $(a,b)$ can be estimated by constrained least squares:
 $\min_{a>0,\;b>0}\;\sum_{i=1}^{n}\bigl[E_{i}-a\,\ln(1+b\,x_{i})\bigr]^{2}.$
-Profiling out $a$ yields, for any fixed $b>0$,
-$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})},\qquad\text{then minimize }S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2}\text{ in }b>0,$ (3)
+For any fixed $b>0$, the optimal $a$ admits the closed-form expression
+$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})}.$ (3)
+Substituting $a(b)$ yields a one-dimensional objective
+$S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2},$
 […]
-If $x$ was fitted in pages with $W$ words/page and you now compute in words, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$; $a$ is unchanged.
+If the model is fitted using pages with $W$ words per page and later applied using word counts, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$, while $a$ remains unchanged.
 7.2 From tolerance to a calibrated score
-For an evaluation of size $x=\text{EWC}$ (words), compute the allowed penalty points at the passing threshold
+For an evaluation sample of size $x=\mathrm{EWC}$ (words), the tolerance threshold at the passing boundary is computed as
 $E_{\text{allowed}}(x)\;=\;\hat{a}\,\ln\!\bigl(1+\hat{b}\,x\bigr).$ (4)
-Let $\mathrm{APT}$ be the absolute penalty total from the MQM table. Define the quality fraction
+Let $\mathrm{APT}$ denote the observed absolute penalty total from the MQM annotation table. We define the quality fraction as
 $\mathrm{QF}(x):=1-\frac{\mathrm{APT}}{E_{\text{allowed}}(x)}.$ (5)
-Map $\mathrm{QF}$ to the standard calibrated scale using Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
+The quality fraction is then mapped onto a calibrated score using the Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
 $\mathrm{OS}(x):=\mathrm{PT}+\mathrm{DPI}\,\mathrm{QF}(x),\qquad\mathrm{OS}_{\text{disp}}:=\min\{\mathrm{MSV},\,\max\{0,\mathrm{OS}(x)\}\}.$ (6)
-The rating is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, else Fail.
+The evaluation outcome is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, and Fail otherwise.
 Monotonicity and clipping.
-Since $\mathrm{OS}(x)=\mathrm{PT}+\mathrm{DPI}\bigl(1-\mathrm{APT}/E_{\mathrm{allowed}}(x)\bigr)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ grows; displaying $\mathrm{OS}_{\rm disp}=\min\{\mathrm{MSV},\max[0,\mathrm{OS}(x)]\}$ does not affect Pass/Fail, which always uses the un‑clipped inequality $\mathrm{APT}\leq E_{\mathrm{allowed}}(x)$.
+Since $\mathrm{OS}(x)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ increases.
+The displayed score $\mathrm{OS}_{\text{disp}}$ is clipped for reporting convenience and does not affect pass/fail decisions, which are determined solely by the inequality $\mathrm{APT}\leq E_{\text{allowed}}(x)$.
 7.3 Scorecard mapping
-We now make the non-linear tolerance usable in everyday QA.
-The table below shows how to compute the scorecard fields from the evaluation word count (EWC), the fitted coefficients $\hat{a},\hat{b}$, and the severity-weighted penalty total (APT).
-The MQM error table itself is unchanged; the only change is that the allowed penalty $E_{\text{allowed}}(x)$ depends on the sample size.
+Table 2 summarizes the computations required to populate an operational scorecard using the evaluation word count (EWC), fitted coefficients $\hat{a},\hat{b}$, and severity-weighted penalty total (APT).
+Importantly, the MQM annotation process and error-type/severity schema remain unchanged; only the tolerance threshold becomes length-dependent via Equation (2).
 Name Excel-style computation
 Evaluation size $x$ (EWC) input
-Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (see §5.3)
-Allowed penalty at size $x$, $E_{\mathrm{allowed}}$ = a*LN(1 + b*EWC)
-Quality fraction QFQF $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
+Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (Section 5.3)
+Allowed penalty $E_{\mathrm{allowed}}(x)$ = a*LN(1 + b*EWC)
+Quality fraction $QF(x)$ $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
 Calibrated score $\mathrm{OS}$ = PT + (MSV-PT) * QF
@@
 Table 2: Scorecard mapping: variables and Excel-style computations.
-This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via Equation  (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
-Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line).
-The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual user tolerance function; the linear line is drawn for comparison only.
-Same-curvature retargeting (optional). If you want a new anchor $E(x_{\text{ref}})=E_{0}$ while maintaining curvature, keep $b$ and solve
+Figure 7 illustrates a real-world non-linear tolerance curve calibrated from institutional anchor points, alongside the corresponding anchored linear rule for comparison.
+Figure 7: Example non-linear tolerance curve (solid) compared with an anchored linear rule (dashed).
+The logarithmic curve is calibrated from tolerance points and represents institutional tolerance as a function of sample size; the linear curve is shown for reference.
+Same-curvature re-anchoring (optional). To impose a new anchor value $E(x_{\text{ref}})=E_{0}$ while preserving curvature, fix $b$ and solve
 $a\;=\;\frac{E_{0}}{\ln(1+b\,x_{\text{ref}})}.$
-This re-anchors the tolerance without changing the curve shape.
 Micro-range caution.
-Below $\sim$250 words, deterministic tolerance curves are statistically unreliable; in this sample size range methods of Statistical Quality Control must be used rather than score curves.
-Quantitatively, if the underlying minor‑error rate is $p$ per word and we observe only $n$ words, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so the relative uncertainty scales as $1/\sqrt{np}$ and becomes huge when $np\approx 1$ (e.g., 1 expected error in 200 words at 5/1000).
+For sample sizes below approximately 250 words, deterministic tolerance curves are statistically unstable, and Statistical Quality Control (SQC) procedures are recommended instead.
+Intuitively, if the underlying minor-error rate is $p$ per word and only $n$ words are observed, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so relative uncertainty scales as $1/\sqrt{np}$ and becomes large when $np\approx 1$.
 Worked example (documentation content).
@@
 For $x=3000$ words, $E_{\mathrm{allowed}}(x)\approx 8.357$ points.
-If $\mathrm{APT}=7$, then $QF=1-7/8.357\approx 0.162$.
+If $\mathrm{APT}=7$, then $QF\approx 0.162$.
 With $\mathrm{PT}=80$, $\mathrm{MSV}=100$ ($\mathrm{DPI}=20$), $\mathrm{OS}\approx 83.25$ and $DM\approx 1.36$ (Pass).
@@
 7.4 Implications for CAT tools and LQA automation
-The change in scoring logic has clear consequences for software systems that implement translation quality evaluation, particularly CAT tools with integrated LQA modules or standalone dedicated LQA platforms.
-In the traditional linear model, these tools calculate scores based on a fixed error density relative to a single reference word count.
-However, with a logarithmic tolerance function, the score calculation must dynamically adjust based on the actual sample size.
-This means LQA modules and dashboards need to:
-  • Replace static pass/fail thresholds based on a single anchor point with dynamic tolerance functions, $E(x)$, that adjust based on the actual word count ($x$).
-  • Visualize quality scores against a curved baseline that represents the tolerance function, not a simple flat line.
-  • Provide evaluators with interactive feedback showing how close a sample’s error count is to its unique, size-adjusted pass/fail threshold.
-This functionality is particularly valuable for machine translation, post-editing, and AI-generated content, as evaluation samples from these sources often vary widely in word count.
-7.5 Improvement of a Unified Calibrated Quality Scale
-One major challenge in TQE is ensuring that quality scores remain human-readable, interpretable, and comparable across different types of content, language pairs, and evaluation sample lengths.
-The linear calibrated model is a significant improvement over using raw scores with very small passing intervals, but even the linear calibrated model can distort comparability when applied to large or small samples.
-With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, user-specific quality scale.
-Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable.
-For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers.
-The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
-This normalization also improves the validity of quality measurement results analytics across a portfolio of projects, enabling better tracking of supplier performance, trend analysis, and root-cause investigations.
+Adopting a logarithmic tolerance function has direct implications for software systems implementing LQA workflows, including CAT tools with integrated quality modules and standalone QA platforms.
+Under proportional scaling, thresholds are defined using a constant error density relative to a fixed reference length; under the proposed model, tolerance must be computed dynamically as a function of the observed evaluation word count.
+In practice, this requires tooling to:
+  • replace static pass/fail thresholds with a length-dependent tolerance function $E(x)$;
+  • visualize quality relative to a curved tolerance baseline rather than a proportional reference line; and
+  • report decision margins (e.g., $DM=E_{\mathrm{allowed}}-\mathrm{APT}$) to communicate proximity to the pass/fail boundary.
+These adaptations are particularly relevant in workflows involving heterogeneous sample sizes, such as machine translation, post-editing, and AI-assisted content generation.
+7.5 Toward a unified calibrated quality scale
+A persistent challenge in analytic TQE is maintaining interpretability and comparability across content types, language pairs, and evaluation sample lengths.
+Calibrated scoring improves over raw penalty rates by mapping results to a stable reporting scale; however, proportional length scaling can still distort comparability when applied outside its effective range.
+By incorporating length-dependent tolerance via $E(x)$, non-linear calibrated scoring supports a unified, institution-specific 0–100 scale for heterogeneous evaluation sizes.
+This improves comparability across jobs and enables more reliable aggregation for analytics such as trend monitoring, supplier benchmarking, and root-cause analysis.
+As a concrete example, [GladkoffEtAl2024] reports a case in which a 5,000-word English–German marketing translation containing 23 minor errors passed under linear scaling but was rejected by expert reviewers.
+The logarithmic model yields a substantially lower tolerance threshold in this regime (16 minor-error equivalents), consistent with the human accept/reject decision.
 7.6 Applications beyond human translation
-While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content.
-Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation [han2024neural].
-The model based on non-linear error tolerance provides a fairer and more realistic baseline for evaluating content.
-It avoids the common pitfalls of unduly penalizing shorter segments with one mistake or overlooking the accumulated minor issues in long, highly fluent AI-generated texts.
-As organizations adopt more automated content generation, the ability to evaluate quality in a human-aligned, perceptually realistic manner becomes essential.
-The proposed scoring model provides a foundation for doing so.
+Although the model is motivated by human translation evaluation practices, the same length-dependent tolerance considerations apply to AI-generated text.
+Large language model (LLM) outputs (e.g., summaries, answers, or translations) are increasingly evaluated using procedures adapted from TQE and LQA [han2024neural].
+Non-linear tolerance modeling provides a more stable baseline for such evaluation by reducing over-penalization of short segments and preventing under-penalization of accumulated minor issues in long outputs.
 7.7 Document-level evaluation and future automation
-Translation quality never exists in a vacuum: every segment lives within the cohesive fabric of a document.
-A translation sample is therefore not a roll of toilet paper to be inspected centimeter by centimeter; its quality must be judged in the context of the larger text.
-The larger the sample relative to the document, the more faithfully that judgment reflects holistic properties such as cohesion, coherence, register, and terminological consistency.
-Small extracts can still provide a rapid signal, but offer only a limited view of document quality.
-They cannot capture discourse-level phenomena, and—because error density interacts non-linearly with length—linear metrics break down outside their narrow reference window.
-The MQM-based Non-Linear Calibrated model solves this by remaining simple and linear in the meso-range yet extending smoothly into the macro-range.
-As long as the evaluation sample is sufficiently large relative to the document, as recommended by the Multi-Range Framework [GladkoffEtAl2024]—it yields scores that are statistically reliable and consistent with the judgment of experts at the document level.
-Of course, analytic TQE becomes increasingly expensive at large evaluation sample sizes.
-No current AI system can yet perform a complete analytic evaluation, but future models may acquire the ability to conduct formal TQE (full analytic MQM evaluation) over long contexts (document-scale samples), reliably capturing local errors (segment-level) and holistic properties (cohesion, coherence, terminology/style consistency, discourse).
-The logarithmic model proposed here is expressly designed to serve as the computational backbone for such future automation, ensuring that automated scoring remains perceptually valid across the full spectrum of document lengths.
+Document-level quality assessment requires integrating local defects with discourse-level properties such as cohesion, register consistency, and terminology management.
+Small samples can provide rapid quality signals, but they provide only partial visibility into document-level phenomena and may lead to unstable conclusions when length effects are ignored.
+The MQM-based non-linear calibrated model addresses this by maintaining linear behavior near its reference regime while extending smoothly to longer samples via a single tolerance function.
+Provided the evaluation sample is sufficiently large to yield statistically stable estimates (as recommended by the Multi-Range framework [GladkoffEtAl2024]), the resulting scores remain consistent with document-level expert judgments.
+Finally, although current automated systems do not yet perform complete MQM-style analytic evaluation over long contexts, future evaluation models may increasingly combine local error detection with document-level assessment.
+In such settings, a length-dependent tolerance function provides a principled backbone for ensuring that automated scoring remains comparable and perceptually aligned across a wide range of evaluation lengths.
 7.8 Summary
-In practical terms, the transition to a non-linear scoring model means:
-  • Greater alignment between future automated document-level scoring systems and expert human judgment,
-  • Improved fairness and consistency across evaluation samples of different lengths,
-  • More intelligent quality evaluation systems producing evaluations closer to human perception,
-  • Stronger foundations for evaluating both human and AI-generated translation.
-This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline.
-The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary.
-It is simple to implement, easy to calibrate, and theoretically grounded.
-By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
+In practical terms, transitioning to a non-linear tolerance model enables:
+  • improved consistency of acceptability decisions across variable-length evaluation samples;
+  • more interpretable and equitable calibrated scores on a unified reporting scale;
+  • simpler integration into CAT and LQA tooling via dynamic tolerance computation; and
+  • better alignment with document-level evaluation requirements for both human and AI-generated content.
+Overall, the proposed approach operationalizes empirically supported non-linear tolerance scaling in a form that is simple to calibrate, straightforward to implement, and compatible with existing MQM-based workflows.
 8 Limitations and Future Work
-The following limitations and research directions qualify the benefits summarized in Section 7.8 and the conclusion.
-While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged.
-Addressing these will be essential for refining the model and validating its broader applicability.
-8.1 Limited Survey Group Size
-We conducted structured interviews with quality managers at three very large enterprise institutes, and the results are corroborated by their internal practices.
-Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small.
-Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy.
-However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
-Further research should involve a broader range of organizations—including LSPs, in-house localization teams, and public-sector translation departments—across different domains and language pairs.
-Ideally, future studies should also include controlled experiments where evaluators independently assess translation quality using both holistic judgment and formal scoring models, enabling direct comparison between model predictions and actual pass/fail outcomes.
-8.2 Content-Type Specific Effects
-Different content types such as marketing, legal, technical, and UI strings, may have different quality expectations and error tolerance.
-Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by user or use case.
-We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters.
-However, this observation requires further validation through broader field data.
-A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend.
-Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms.
-We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
-8.3 Severity Weighting and Mixed Error Types
-This paper concentrates on modeling tolerance in terms of penalty points, each of which typically represents one minor error.
-In practice, quality evaluations combine several error severities (minor, major, critical) penalized by specific multipliers and sometimes further scaled by different error-type weights applied for different content types.
-Severity scaling, i.e., fixed severity weights such as minor/major/critical multipliers, is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points.
-This is separate from any error-type weights (e.g., terminology vs.
-grammar) a metric may also define.
-While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often.
-For these rarer severities, an integer-valued sequence may be preferable, as indicated by Institution 3’s data in Fig.
-3.
-That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
-Future work should examine the form of such integer sequences, quantify how far they depart from a pure logarithmic curve, and determine the parameters that best describe severity scaling across different content types and languages.
-8.4 Inter-Rater Reliability, IRR
-When a linear model is used, the variability among human raters can be partly attributed to the fact that the model is only accurate at its anchoring point and loses precision elsewhere.
-In contrast, a calibrated non-linear model better reflects human judgment across a wide range of evaluation sample sizes.
-As a result, Inter-Rater Reliability (IRR) should increase significantly when using a non-linear calibrated scoring model, as it provides consistent and comparable measurements across different evaluation sample sizes, languages, and content types.
-More research is needed to quantify the improvement in inter-annotator agreement and to determine how much such a model reduces variance compared to traditional linear approaches in real-world annotation settings.
+This section qualifies the practical benefits summarized in Section 7.8 and identifies directions for further validation and extension.
+Although the proposed non-linear tolerance model is supported by both expert elicitation and theoretical motivation, several limitations constrain the scope of the current evidence base.
+8.1 Sample size and elicitation design
+The empirical evidence in Section 3 is based on structured elicitation with Quality Managers from three large institutional translation programs.
+While these organizations operate mature, high-volume multilingual workflows, the number of participating institutions (and respondents) remains limited.
+Moreover, tolerance judgments were elicited rather than measured in blinded controlled experiments, which may introduce subjectivity and potential response bias.
+Future work should expand the sample to include a broader range of stakeholders—e.g., LSPs, in-house localization teams, and public-sector translation units—across domains and language pairs.
+Controlled studies could further strengthen validity by comparing (i) holistic accept/reject decisions and (ii) analytic MQM-based scoring outcomes under alternative tolerance functions, enabling direct assessment of predictive alignment and decision consistency.
+8.2 Content-type and domain effects
+Quality expectations vary by content type (e.g., marketing, legal, technical documentation, UI strings).
+The proposed model assumes a common functional form $E(x)=a\ln(1+bx)$, with content- or user-specific variation captured by the parameters $a$ (scale) and $b$ (curvature).
+Our operational observations suggest that a single logarithmic form can perform robustly across heterogeneous content types, but broader field evidence is required.
+An important extension is to test whether particular domains exhibit systematically different curvature profiles (e.g., steeper for high-risk instructional content or flatter for low-stakes internal communication), and whether additional saturation occurs for very large evaluation sizes (e.g., book- or manual-length contexts).
+If such effects are observed, alternative sublinear families (e.g., log–log, logistic saturation, or piecewise concave forms) may be warranted.
+8.3 Severity weighting and mixed error distributions
+This paper models tolerance in severity-weighted penalty points, where one point is typically interpreted as one minor-error equivalent.
+In operational settings, evaluations include mixed severities (minor/major/critical) with fixed multipliers and may also include error-type weights.
+Although the logarithmic tolerance curve captures minor-error tolerance well—minor errors being the most frequent and thus statistically most stable—rarer severities may require discrete (integer) decision rules or severity-specific tolerance sequences.
+Future work should evaluate how tolerance behaves for major and critical errors under variable-length sampling.
+This includes establishing principled rounding and decision rules, testing whether severity-specific tolerance sequences remain approximately sublinear, and assessing whether severity multipliers interact with length-dependent tolerance in ways that call for multi-parameter extensions.
+8.4 Inter-rater reliability
+A key motivation for length-aware tolerance is to reduce systematic disagreement caused by applying proportional extrapolation outside its effective reference region.
+The proposed model is expected to improve inter-rater reliability (IRR) by providing more consistent pass/fail thresholds across varying sample sizes.
+However, the present work does not yet quantify IRR improvements under controlled rater studies.
+Future work should measure IRR (e.g., using Krippendorff’s $\alpha$, Cohen’s $\kappa$, or intra-class correlation, depending on task design) under both linear and non-linear scoring regimes.
+Such studies could test whether a length-dependent tolerance model reduces variance in borderline cases and improves agreement across evaluators with different experience levels.
 8.5 Extension to AI evaluation benchmarks
-This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models.
-However, this claim has yet to be formally validated.
-AI benchmarks such as WMT [^10] https://machinetranslate.org/wmt shared task series [kocmi-etal-2025-findings] which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
-Future work should involve testing the non-linear model as a normalization layer in LLM evaluation tasks, particularly those involving paragraph- or document-level generation, where holistic perception plays a major role.
-A further implication is that a calibrated non-linear scoring model could eventually serve as a reference framework for future AI-based document-level evaluations, which currently provide holistic but qualitative judgments without a quantitative calibration.
+We hypothesize that similar length-dependent tolerance effects may apply to the evaluation of AI-generated text (including document translation, summarization, and dialogue).
+This claim remains to be validated empirically.
+Benchmarking settings—including shared-task evaluations in MT such as WMT [kocmi-etal-2025-findings]—often rely on reference-based metrics or length-normalized error densities; incorporating calibrated non-linear normalization may improve comparability across variable-length outputs and strengthen alignment with human preferences in document-level evaluation.
+Future work should test the proposed model as a normalization layer for paragraph- and document-level evaluation in both MT and broader LLM generation tasks, including analysis of robustness across domains and prompt conditions.
+A longer-term direction is to investigate whether calibrated tolerance functions can serve as quantitative scaffolding for AI-assisted document-level quality assessment systems that currently produce holistic but weakly calibrated judgments.
 9 Conclusion
-Translation quality evaluation plays a critical role in ensuring that multilingual content meets the linguistic, technical, and functional standards required for global communication.
-While current industry practices predominantly rely on linear scoring models—extrapolating error thresholds proportionally from fixed reference wordcounts—this paper builds on empirical and practitioner evidence showing that linear models do not consistently align with human evaluation, motivating the adoption of non-linear alternatives.
-This paper presents empirical data from experienced quality managers at large institutional translation service users, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size.
-This effect was evident across organizations, content types, and error categories and severities.
-This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
-To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance.
-This model is mathematically simple, empirically grounded, and easily adaptable to specific user requirements.
-It provides a flexible, yet rigorous alternative to traditional linear error density calculations.
-Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
-The implications of adopting this model are wide-ranging.
-Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths.
-CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics.
-Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output.
-While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear.
-By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
-Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable.
-In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method.
-[^11] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes.
-SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output).
-In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making.
-This study consolidates the multi-range theory findings by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
+This paper addressed the length-dependence of acceptability thresholds in MQM-based analytic Translation Quality Evaluation (TQE).
+While proportional (linear) error scaling is widely used in practice, our results indicate that it can produce systematic bias when evaluation sample sizes deviate from the reference word count.
+First, we provided empirical evidence from three large institutional translation programs showing that tolerable error counts increase with sample length sublinearly, yielding concave tolerance profiles consistent with logarithmic growth.
+Second, we linked this behavior to established theoretical accounts from psychophysics (compressive perceptual scaling) and Cognitive Load Theory (accumulated processing disruption), which together motivate non-linear integration of errors over increasing text length.
+Third, we proposed a calibrated two-parameter tolerance function $E(x)=a\ln(1+bx)$ and showed how it can be anchored from tolerance points using a simple one-dimensional calibration procedure.
+Fourth, we demonstrated that the model can be integrated into existing MQM pipelines with minimal change: the MQM annotation process and penalty computation remain unchanged, while a constant length-normalized tolerance is replaced by a length-dependent threshold function.
+In addition, we derived a near-reference fidelity interval that characterizes when a linear approximation remains sufficiently accurate (e.g., within $\pm 20\%$), providing a practical guideline for model selection in operational deployments.
+Finally, we reiterate that for very short samples (approximately $\lesssim 250$ words), deterministic tolerance curves are statistically unstable, and risk-based Statistical Quality Control (SQC) remains the appropriate framework for decision-making under high sampling uncertainty [GladkoffEtAl2024].
 10 acknowledgments
@@
 Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
-Appendix B Linear vs Logarithmic Fits
-For readers who want the formal evidence behind Fig.
-3: we fit both a straight line and a logarithmic curve to the same elicited points and compared their goodness-of-fit using standard metrics.
-Smaller is better for SSE (sum of squared errors) and RMSE (root-mean-square error); larger is better for $R^{2}$ (coefficient of determination).
-AIC (Akaike Information Criterion) and BIC (Bayesian Information Criterion) are model-selection scores that penalize extra parameters; lower is better.
+Appendix B Linear vs. Logarithmic Fits
+To complement the qualitative trend in Figures 13, we compare linear and logarithmic functional forms by fitting multiple candidate models to the same elicited tolerance points and evaluating goodness-of-fit.
+We report standard error-based measures (SSE, RMSE), explained variance ($R^{2}$), and information criteria (AIC, BIC), where lower values indicate better fit for SSE/RMSE/AIC/BIC and higher values indicate better fit for $R^{2}$.
+Fit metrics.
+For observed points $\{(x_{i},y_{i})\}_{i=1}^{n}$ and fitted values $\hat{y}_{i}$, the metrics are defined as:
 $\mathrm{SSE}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\qquad\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
-$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
-where $n$ is the number of points and $k$ the number of fitted parameters.
-As Table 3 shows, the logarithmic model has dramatically smaller SSE/RMSE, much higher $R^{2}$, and much lower AIC/BIC, so its advantage is not an artifact of the extra parameter but a closer match to the elicited judgments.
-These all confirm, quantitatively, that the log model explains about $95\%$ of the variance ($R^{2}\!=\!0.94$), while the linear model explains very little ($R^{2}\!=\!0.04$), and that the log model’s advantage is not merely due to having one extra parameter.
-In other words, the visual impression is also statistically decisive.
-Below is the quantitative comparison of linear vs.
-logarithmic fits on the elicited points.
-To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes.
-This is what the values in Table 3 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better).
-Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC.
-This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can.
-In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
-Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
+$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}},\qquad\mathrm{AIC}=n\ln(\mathrm{SSE}/n)+2k,\quad\mathrm{BIC}=n\ln(\mathrm{SSE}/n)+k\ln n,$
+where $n$ is the number of points and $k$ is the number of fitted parameters.
+Models.
+We compare three candidate forms:
+(i) Logarithmic: $y=a\ln(1+bx),$
+(ii) Linear (with intercept): $y=\alpha+\beta x,$
+(iii) Linear (through origin): $y=cx.$
+Model (iii) is included because proportional tolerance scaling implicitly assumes a zero intercept (i.e., no tolerance at zero length), whereas Model (ii) provides a standard unconstrained linear baseline.
+Table 3 reports fitted parameters and fit statistics for the Institution 1 minor-error elicitation dataset (data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)).
+Model $a$ $b$ $\alpha$ $\beta$ SSE RMSE $R^{2}$ AIC BIC
 Logarithmic $a\ln(1+bx)$ $3.353$ $0.590$ $1.551$ $0.471$ $0.945$ 6.550-6.550 6.658-6.658
-Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
-AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7, k=2k{=}2 (log) and k=1k{=}1 (linear). Smaller is better.
-Table 3: Goodness-of-fit statistics for the user 1 minor-error dataset ($x$ in pages; $E$ in minor points).
-Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit.
-Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
-The standard indicators reported in Table 3 show how well a curve follows the observed data points.
-They illustrate why the logarithmic model offers a markedly superior fit than the linear one.
-[^12] SSE (sum of squared errors): the total squared distance between the curve and the data (smaller is better).
-RMSE: the average size of that distance (smaller means the model sits almost on top of the data).
-$R^{2}$: the fraction of variability explained by the curve ($R^{2}\approx 1$ means excellent, $R^{2}\approx 0$ means poor).
-AIC and BIC: compact comparison measures that reward better fit but penalize extra parameters (lower is better).
-Thus, even without going into formulas, the table makes visible that the non-linear model is overwhelmingly better at reflecting real human tolerance.
-When multiple elicited points are available, we recommend weighted least squares (App.
-B), down‑weighting sparse or less reliable page sizes.
-Using the delta method in App.
-B, we obtain approximate prediction standard errors and plot shaded confidence bands around the fitted curve.
+Linear (with intercept) $\alpha+\beta x$
+Linear (through origin) cxcx $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
+AIC $n\ln(\mathrm{SSE}/n)+2k$, BIC $n\ln(\mathrm{SSE}/n)+k\ln n$; here n=7n{=}7. Parameter counts: k=2k{=}2 for logarithmic, k=2k{=}2 for linear with intercept, and k=1k{=}1 for linear through origin.
+Table 3: Goodness-of-fit statistics for Institution 1 minor-error elicitation ($x$ in pages; $y$ in minor-error equivalents).
+Comparison.
+The logarithmic model substantially improves fit compared to a purely proportional linear rule.
+Reporting both constrained and unconstrained linear baselines strengthens the robustness of the comparison and avoids conflating functional form with parameter constraints.
 Appendix C Calibration from Two Points (Numerical)